### v1

# Inattention to States and Characteristics

Abstract
We study a generalization of the state-action rational inattention model with two dimensions of uncertainty: states and hedonic characteristics. The resulting conditional choice probability takes a familiar weighted multinomial logit form, but in contrast to the state-action model, is unique and is not a corner solution. The model imposes testable restrictions on the behavior of choice probabilities in markets with product entry. The optimal information policy can be found by optimizing subject to a marginal constraint, then finding the optimal marginal. The solution to the former is the Schrödinger bridge of an entropic optimal transport problem. The latter is solved by maximizing a strictly concave function (over a convex set) which envelops the objective.
Introduction
Consumer choice typically contains two dimensions of uncertainty: one concerns the product, and the second concerns utility-relevant future states of the world. When buying insurance, for example, the consumer faces uncertainty surrounding the contents of each insurance contract and their own future risk.
An increasingly popular framework for studying decision-making under uncertainty is the state-action rational inattention model [4]. In it, the decision-maker (DM) commits to an information policy consisting of a joint distribution $P$ over actions and states, subject to a Bayes plausibility constraint. The DM seeks to maximize expected utility less an information cost which is linear in the Shannon information between state and action. A relationship between the optimal conditional choice probability and the optimal marginal probability is known; however, the optimal marginal is not necessarily unique, and therefore neither is the optimal conditional.
We present a natural extension of the state-action model, which we call the state-characteristic model. In this model, the DM faces an exogenous prior distribution $\phi$ over characteristics $\xi$ – reflecting the inherent ubiquity of certain characteristics – and commits to a joint distribution $P$ over characteristics $\xi$ and states $\theta$. Mathematically, our model differs in only one respect: in addition to the usual Shannon information cost, we add a cost of diverging from $\phi$, interpretable as information gain from $P$ over $\phi$. This models, for instance, the fact that in a market full of lemons, one would need to acquire information to ensure that the likelihood of buying lemons is low.
The dual information costs account for the two dimensions of information gain that is observed by the econometrician when they observe the joint distribution over $(\xi,\theta)$. First, the econometrician observes there is information gain from $P(\xi|\theta)$ over $\phi(\xi)$ – that is, the choice distribution over characteristics is not what it would be if consumers chose at random. Second, the econometrician observes that the consumer’s choice of coverage level provides information about what their future losses will be – which differs from the prior $\mu$.
In section˜2, we motivate the need for the state-characteristic model by showing that the state-action model cannot account for both dimensions of information observed by the econometrician: modeling characteristics as actions fails to account for information gain from $P(\xi|\theta)$ over $\phi(\xi)$, while modeling characteristics as stochastic fails to account for information gain from $P(\theta|\xi)$ over $\mu(\theta)$. The state-characteristic model can be thought of as an interpolation between these two approaches which includes both information costs. We present the model in full generality in section˜3.
In our model, the relative costs for each dimension of information is parametrized by $\alpha\in[0,1]$. Our model thus nests the two aforementioned versions of the state-action model discussed in section˜2 at $\alpha=0$ and $\alpha=1$. As results have already been established for these cases, we focus on the case $\alpha\in(0,1)$.
Our main result, presented in section˜4, is very similar to the famous weighted multinomial logit (MNL) result [4]. Whereas their optimal conditional choice probability (CCP) satisfied
$P(a|\theta)=\frac{P(a)e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
for some normalizing partition function $Z$, our CCP satisfies
$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
One major advantage that our model ($\alpha\in(0,1)$) has is it always admits a unique solution whereas state-action ($\alpha=0$) does not. Therefore, even in cases where state-action seems to better model the situation, the state-characteristic model may nevertheless be useful for selecting from multiple solutions. Another distinction is that in state-action model, the DM forms (possibly non-unique) optimal “consideration sets” [1] – such that actions outside the set have $P(a)=0$ – whereas a key feature of the state-characteristic model is that no characteristic is completely ruled out.
This reflects the idea that consumers typically content themselves with being pretty sure of certain things; being truly 100% sure of anything requires an irrational amount of effort. For example, to acquire information as to how sausages are made, the conscious consumer may read the nutrition facts, the concerned consumer might read the ingredients, but nobody burglarizes a sausage factory with a spectrometer to make sure that the sausages are not produced in an environment with trace amounts of soy.
In lay terms, what the MNL formula states is that the CCP depends on three things. The first is the utility. The second is the inherent commonality of characteristics $\phi$: the more lemons there are, the more likely it is that you will buy one. The third is the marginal probability $P$. If you prefer $\xi$ to $\xi^{\prime}$ in 99% of states, then even in the 1% of states where you slightly prefer $\xi^{\prime}$ to $\xi$, you may still end up picking $\xi$ unless you acquire enough information about the state to be quite sure you are in that 1%.
How much each of these three things matters depends on the parameters $\alpha,\lambda$. If $\alpha\lambda$ is large, that means it is harder to learn about the characteristics $\xi$. Thus, $P(\xi|\theta)$ will resemble what would happen if you were to choose randomly, which is represented by the prior $\phi(\xi)$. If $(1-\alpha)\lambda$ is large, that means it is harder to learn about states $\theta$. Thus, $P(\xi|\theta)$ will not differ much from $P(\xi)$, since separating $P(\xi|\theta)$ from $P(\xi)$ means you learned to differentiate $\theta$ from other states. If both are large, then $P(\xi)$ is close to $\phi(\xi)$ and the CCPs are close to both. Conversely, as $\lambda\to 0$, the CCP begins to look like that of a fully-informed, utility (soft-)maximizing agent.
Lastly, section˜5 concerns falsifiable restrictions implied by the results. In particular, entry of products into the market which generate variation in $\phi$ should generate changes in CCP ratios that are not state-dependent. These predictions do not require knowledge of the prior $\phi$ to be tested. However, if the prior is known, the relative cost parameter $\alpha$ is over-identified in the process.
In the finite discrete case, the results may be established by taking Lagrangians. In the remaining sections, we derive the results for the general case, since characteristics may be continuous. The results hereafter are mostly mathematical in nature, whose utility lies in that they provide (hopefully) interesting and (hopefully) novel insights into the problem, as well as connections to other well-known problems. The proofs in these sections are typically deferred to the appendix.
Section˜6 explores the first-step orthogonality property of Shannon entropy-based inattention objectives and its consequences. In brief, the objective (which is a function of the information policy $P$) can be written as an integral in which $P$ appears twice: once in the integrand and once in the form of $dP$. While the product rule suggests that perturbing $P$ ought to affect the objective in two ways, first-step orthogonality establishes that it only locally affects the objective in one: via the re-weighting in the form of $dP$. First-step orthogonality implies that both state-action and state-characteristic have a nice property whereby when the objective is written as a single integral, the integrand is constant.
Section˜7 establishes a connection between rational inattention and entropic optimal transport (EOT). It turns out that the problem of maximizing the state-characteristic objective subject to a second marginal constraint $P|_{\xi}=\nu$ (on top of Bayes plausibility, which is itself a marginal constraint $P|_{\theta}=\mu$) can be re-cast as a Schrödinger bridge problem [7]. The problem comes with a dual which consists of finding Schrödinger potentials $a(\xi),b(\theta)$ which satisfy the associated Schrödinger equations. From EOT theory, we know that the value as a function of the marginal constraint $\nu$ is additively separable and equal to the sum of (the expectations of) the potentials. The potentials have a straightforward Lagrangian interpretation. Loosely speaking, the potential associated with the characteristic $\xi$ corresponds to how much the value would change if we were to change the marginal constraint $\nu$ such that some extra mass were added to $\xi$.
Perhaps the most interesting practical takeaway here is that the optimal CCP subject to two marginal constraints is also a weighted MNL. [4] had established that
$\text{$P$ is optimal}\implies P(\xi|\theta)\propto P(\xi)e^{u(\xi,\theta)/\lambda}$
but this does not necessarily imply that
$\text{$P$ is optimal subject to }P|_{\xi}=\nu\implies P(\xi|\theta)\propto\nu(\xi)e^{u(\xi,\theta)/\lambda}$
Indeed, it is not true. However, the Schrödinger equations imply that $P$ maximizes the objective subject to $P|_{\xi}=\nu,P|_{\theta}=\mu$ if and only if there exists $a,b$ such that
$P(\xi|\theta)=\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}e^{[u(\xi,\theta)-a(\xi)-b(\theta)]/\lambda}$
Thus, $P(\xi|\theta)$ always follows a weighted MNL. Finding such an $a,b$ when $\xi,\theta$ are discrete is a case of the matrix scaling problem. Sinkhorn’s algorithm can be used to find the Schrödinger potentials.
We finish by considering the problem of finding the optimal marginal in section˜8. In [1], the authors write out a concave function over the set of marginals which helps them establish that the Kuhn-Tucker conditions are sufficient, as opposed to only necessary. This function may be obtained by swapping an expectation with a log; by Jensen’s inequality, it upper bounds the objective. It turns out that this function has even nicer properties when $\alpha\in(0,1)$. In particular, it is strictly concave over a convex set and coincides with the objective at their mutual optima. Thus, rather than maximize a more complicated objective, it suffices to solve this much easier convex optimization problem.
Motivation
State-action presents two options for modeling consumer choice with uncertain characteristics: either treat the characteristics as actions – i.e. costlessly observed – or as a state variable, alongside other state variables. The latter option, however, presents an identification problem. It turns out that a DM who does not observe the other state variables can achieve the same payoff at the same information cost as someone who costlessly observes them. Thus, the two behave identically and are observationally equivalent.
Example. Consumers choose between two insurance contracts $j=1,2$. They know one is high-coverage $\xi_{j}=H$ and the other is low-coverage $\xi_{j}=L$ but do not know which is which. The consumer has also two future health expenditure states, $t\in\{H,L\}$, each with probability $1/2$. The consumer receives utility 1 if they choose the ‘correct’ contract (high coverage iff high future expenditures) and 0 otherwise. In other words,
$u(j,\xi_{1},\xi_{2},t)=\textbf{1}\{\xi_{j}=t\}=\begin{cases}1&\xi_{j}=t\\ 0&\xi_{j}\neq t\end{cases}$
Recall the state-action sequence proceeds as follows: the consumer commits to an information policy – a Bayes plausible joint distribution over actions $j$ and states $(\xi_{1},\xi_{2},t)$. Nature then draws an action recommendation and a state. Now, suppose consumer 1 observes $t$ before committing to the information policy
if $\xi_{1}=t$, recommend $j=1$ with probability $p$ and $j=2$ with probability 1p1-p
if $\xi_{2}=t$, recommend $j=2$ with probability $p$ and $j=1$ with probability 1p1-p
and consumer 2 does not observe $t$, but commits to the same information policy. Consider:
  1. 1.
    Do the two consumers have the same information upon receiving the same recommendation $j$?
  2. 2.
    Do the two consumers pay the same information cost?
  3. 3.
    Is the expected payoff (averaged over $t$) the same?
The answer to the first, clearly, is no. Consumer 1 knows $t$ and $j$, whereas consumer 2 only gets $j$. Consumer 1 knows, with probability $p$, that the coverage level will be $t$, whereas consumer 2 merely knows that with probability $p$, the coverage level will be appropriate for their future health expenditures – but does not know what either are. The answer to the second is yes. The answer to the third is also yes – for both, their utility will be 1 with probability $p$ and 0 with 1p1-p.
Ergo, consumer 2 is able to achieve the same expected payoff at the same information cost, despite having less information before committing to the information policy and after receiving the recommendation. Worse, for the econometrician, the conditional choice probability for both consumers is the same, so they are indistinguishable.
What is happening is that both consumers are acquiring the same amount of information about whether $j$ is the “right” choice. Consumer 1 already knows $t$ and is doing so by acquiring information about what $j$ is. But consumer 2 is doing so without acquiring any information about what $j$ is or why it might be the right choice. What the state-action model accounts for is the fact that for each $t$, the probability of choosing $\xi_{j}=t$ differs from $1/2$ – what it would be if no information was acquired; what it doesn’t account for – or rather, treats as costless – is the information about $t$ itself. This issue can be formalized as follows.
Observational equivalence in state-action models when characteristics are state variables. In general, could an econometrician determine whether an insurance buyer in the state-action model has a priori private information about their future health expenditures $\theta$? Consider the canonical consumer problem setting (see e.g. [1]), the action space is $A=\{1,...,J\}$, and the utility vector is $\textbf{u}(\boldsymbol{\xi},\theta)=(u(1,\boldsymbol{\xi},\theta),...,u(J,\boldsymbol{\xi},\theta))$, which is a function of the random variables $j\in A$, $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$, and $\theta\in\Theta$. The state-action model predicts that the optimal conditional choice probability is
$P(j|\boldsymbol{\xi},\theta)=P(j|\textbf{u}(\boldsymbol{\xi},\theta))=\frac{e^{u(j,\boldsymbol{\xi},\theta)}}{\sum_{k\in[J]}e^{u(k,\boldsymbol{\xi},\theta)}}$
This equation holds regardless of whether $\theta$ is assumed to be known to the consumer a priori or not. Thus, whether $\theta$ is known a priori is not identified in the state-action model.
Mathematically, we observe that by the law of iterated expectations,
$I_{P}(j,\textbf{u}):=\int\log\left(\frac{dP(j,\textbf{u})}{dP(j)dP(\textbf{u)}}\right)\>dP=\iint\log\left(\frac{dP(j,\textbf{u}|\theta)}{dP(j|\theta)dP(\textbf{u}|\theta)}\right)\>dP(j,\textbf{u}|\theta)\>dP(\theta)=\int I_{P}(j|_{\theta},\textbf{u}|_{\theta})\>dP(\theta)$
because $P(j|\theta)=P(j)$. So, the DM maximizes
$E_{P}[u]-\lambda I_{P}(j,\textbf{u})\equiv E_{P}\bigg{[}E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})\bigg{]}$
if and only if they maximize $E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})$ for almost every $\theta$. In words, the DM solves the state-action rational inattention problem overall if and only if they solve the state-action rational inattention problem in every health state $\theta$.
One final observation is that $I_{P}(j|_{\theta},\textbf{u}|_{\theta})=D_{KL}(P(\xi|\theta)\|\phi(\xi))$ for a natural prior $\phi$. To see this, suppose that out of the $J$ options, $k_{x}$ of them had the characteristic $x\in X$. A natural prior, then, would be $\phi(x)=k_{x}/J$. Without loss, suppose $\xi_{j}=x$. Then, $P(\xi_{j}|\theta,\boldsymbol{\xi})=P(x|\theta)/k_{x}$. Thus,
$\frac{P(j|\textbf{u})}{P(j)}=\frac{P(j|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(\xi_{j}|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(x|\theta)/k_{x}}{1/J}=\frac{P(x|\theta)}{\phi(x)}$
The need for consumer choice uncertainty. Given the non-identifiability problem, it may be tempting to simply assume that a consumer could observe the characteristics of their choices costlessly, and simply use the state-action model to model the uncertainty generated by $\theta$. Indeed, this is what most empirical discrete choice models implicitly do when $\xi$ is taken to be the choice variable. This is reasonable since the econometrician observes $P(\xi,\theta)$, not $P(k,\textbf{u})$. However, there are many circumstances in which the state-characteristic model we propose herein may be helpful.
For one, the choice probability predicted by the state-action model may be non-unique. Second, we may want to incorporate information that some options are more likely than others. Third, we may observe choices which appear strictly dominated, as is empirically common [2], which cannot be explained by modeling characteristics as actions in the state-action model. Fourth, there is no incorporation of the inherent ubiquity of certain options.
To illustrate these problems in structural econometric context, suppose Harold Zurcher is buying used buses, and $X=\{\text{red bus, blue bus, lemon bus}\}$ such that $\text{red}=\text{blue}\succ\text{lemon}$ in all states. In the state-action model without choice uncertainty, he would ever choose a lemon, no matter how abundant they were. And even if we knew red buses outnumbered blue buses 100-to-1, the state-action model does not prescribe a unique optimum policy – it could be optimal for Zurcher to only buy blue (or red).
So what are we missing? What we have shown is that when characteristics are modeled as state variables, the state-action imposes a cost function which is effectively $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$, and when characteristics are modeled as actions, then naturally the cost function is $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$. In the former, the DM behaves as though observing $\theta$ is free. In the latter, the DM observes $\xi$ for free.
To wit, the natural extension of the state-action model which we propose in the next section contains two terms: $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$ and $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$.
Model
Notation. Let $m,m^{\prime}$ be measures defined on the same measurable space. $m\ll m^{\prime}$ denotes absolute continuity. $m\ll m^{\prime}\ll m$ is interchangeable with $\sim m^{\prime}. m=m$$m^{\prime}$ if $m(A)=m^{\prime}(A)$ for each measurable set. The KL-divergence between two probabilities is defined
$D_{KL}(m\|m^{\prime})=\begin{cases}\int\log\left(\frac{dm}{dm^{\prime}}\right)\>dm&(m\ll m^{\prime})\\ \infty&\text{else}\end{cases}$
By convention, $\infty\cdot 0=0$ for integration purposes.222see [3]
Model. Fix an ambient probability space $(\Omega,{\cal F},\mathbb{P})$, a space of characteristics $X$ with an exogenous prior probability $\phi\in\Delta X$, and a space of states $\Theta$ with prior probability $\mu\in\Delta\Theta$. The decision-maker’s (DM) problem is to maximize expected utility $u$ less an information cost $\kappa_{\alpha,\lambda}$
$\sup_{P\in\Delta(X\times\Theta)}U(P):=E_{P}[u(\xi,\theta)]-\kappa_{\alpha,\lambda}(P)$ (1)
subject to
  1. 1.
    (Bayes plausibility) $P|_{\theta}=\mu$
  2. 2.
    (marginal feasibility) $P|_{\xi}\ll\phi$
The cost function $\kappa_{\alpha,\lambda}$ can be written in several ways:
[…]
$\equiv\alpha\lambda\bigg{(}\underbrace{D_{KL}(P(\xi)\|\phi(\xi))+I_{P}(\xi,\theta)}_{\text{information about }\xi}\bigg{)}+(1-\alpha)\lambda\underbrace{I_{P}(\xi,\theta)}_{\text{information about $\theta$}}$
$\equiv\alpha\lambda\underbrace{D_{KL}(P(\xi)\|\phi(\xi))}_{\text{``vertical information''}}+\lambda\underbrace{I_{P}(\xi,\theta)}_{\text{``horizontal information''}}$
where $I_{P}$ denotes mutual information under $P$. To generalize, we impose further regularity conditions:
  1. 1.
    $X,\Theta$ are Polish spaces with Borel $\sigma$-fields ${\cal B}_{X},{\cal B}_{\Theta}$, over which $\phi$ and $\mu$ are respectively defined
  2. 2.
    $u:X\times\Theta\to[\underline{u},\overline{u}]$ is real-valued, bounded, and measurable
Example: state-action with costlessly observed $\xi$. When $\alpha=0$, $\kappa_{\alpha,\lambda}(P)$ simplifies to $I_{P}(\xi,\theta)$, which corresponds to the state-action rational inattention model. Id est, when $\alpha=0$, the DM can choose characteristics directly, just as in the state-action model, the DM chooses actions directly.
Example: state-action with costlessly observed $\theta$. Consider a state-action model in which the DM knows their type $\theta$ a priori. The DM has an action set $A=\{1,...,J\}$ consisting of $J$ products with $J$ characteristics. The DM knows what the characteristics are, but does not know how they are permuted among the $J$ actions. Let $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$ denote these characteristics. Denote $\xi$ with the characteristic of the product he eventually chooses. The mutual information $I_{P}(j,\boldsymbol{\xi})$ between action $j\in[J]$ and $\boldsymbol{\xi}$ is
$\int D_{KL}(P(\boldsymbol{\xi}|j)\|P(\boldsymbol{\xi}))\>dP(j)=D_{KL}(P(\boldsymbol{\xi}|j=1)\|P(\boldsymbol{\xi}))=D_{KL}(P(\xi_{1}|j=1)\|P(\xi_{1}))=D_{KL}(P(\xi)\|\phi(\xi))$
The third equality follows from the observation that choosing $j=1$ provides only information about what $\xi_{1}$ and is not (meaning it provides no information about $(\xi_{2},...,\xi_{J})$ beyond the fact that they are not $\xi_{1}$). Thus, this corresponds to the state-characteristic model with $\alpha=1$ and $\phi(\xi)=1/J$.
Example: Maxwell-Boltzmann statistics. When $\alpha=1$, the law of iterated expectations yields
$U(P)$ $=E_{\theta\sim\mu}\bigg{[}E_{P}\bigg{[}u(\xi,\theta)-\lambda D_{KL}(P(\xi|\theta)\|\phi(\xi))\ |\ \theta\bigg{]}\bigg{]}$
$=E_{\theta\sim\mu}\bigg{[}E_{P}[u(\xi,\theta)|\theta]-\lambda\cdot E_{\phi}\bigg{[}\frac{dP(\xi|\theta)}{d\phi(\xi)}\log\left(\frac{dP(\xi|\theta)}{d\phi(\xi)}\right)\bigg{]}\bigg{]}$
That is, $U(P)$ is maximized iff the conditional distribution maximizes what is inside the expectation for every $\theta$. Independent of the forthcoming results, the Donsker-Varadhan variational formula implies the conditional choice probability follows Maxwell-Boltzmann statistics:
$P(\xi|\theta)=\frac{\phi(\xi)e^{u(\xi,\theta)/\lambda}}{Z(\theta;P)}$
From a physical view, the utility maximization problem can be formulated as a free energy minimization problem. The prior $\phi$ corresponds to the “degeneracy” of the energy level. The utility corresponds to the enthalpy. $\lambda$ corresponds to the temperature. $E_{\phi}\bigg{[}\frac{dP(\xi|\theta)}{d\phi(\xi)}\log\left(\frac{dP(\xi|\theta)}{d\phi(\xi)}\right)\bigg{]}$ is the conditional differential entropy (the cross-entropy between $P(\cdot|\theta)$ and $\phi$).
Measure-theoretic notes. Some care must be taken when considering the distinction between a function $f$ and its associated equivalence class of functions $g$ for which $f=g$ a.e. In general, the ‘choice’ of $P$ does not precisely pin down its disintegration kernels or associated densities. Though not fatal, the ambiguity of possibly infinite densities clutters analysis with qualifiers. To simplify things, we restrict the DM to choosing disintegrations and densities which satisfy
$\frac{dP(\xi,\theta)}{d(P|_{\xi}\otimes\mu)(\xi,\theta)}=\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}<\infty$ $\frac{dP(\xi)}{d\phi(\xi)}<\infty$ $\frac{dP(\theta)}{d\mu(\theta)}=1$ (2)
everywhere, not just almost everywhere. The existence of these are established in the following lemmas.
Lemma 1.
If $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$, then $P(\cdot|\theta)\ll\phi$ $\mu$-a.s. and $P\ll\phi\otimes\mu$.
Lemma 2.
Let $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$. There exist valid densities such that eq.˜2 is satisfied.
The proofs are deferred to the appendix.
Main Results
We consider the case where $\alpha\in(0,1)$. We normalize $\lambda$ to $1$, since maximizing $U(P)$ is the same as maximizing $U(P)/\lambda$. To obtain results for $\lambda\neq 1$, simply take all results herein in which $u(\xi,\theta)$ and associated additive terms appear and divide them by $\lambda$.
Theorem 3 (Gibbs property).
If $P$ solves the objective eq.˜1 then
$\xi\mapsto Y(\xi,\theta;P)$ is constant $P(\cdot|\theta)$-a.s. for $\mu$-a.e. $\theta$
Corollary 3.1 (Weighted multinomial logit).
If $P$ solves the objective eq.˜1 then
$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(\xi,\theta)}}{Z(\theta;P)}$ (3)
for some normalizing partition function $Z$.
Theorem 4.
The optimal marginal $P|_{\xi}$ is unique, and therefore the optimal CCP is unique.
Theorem 5.
If $P$ maximizes $V$, then $P|_{\xi}\ll\phi\ll P|_{\xi}$.
These follow from corollaries˜7.1, 8 and 15.1. In the discrete case, these can be derived by taking Lagrangians.
Testable restrictions
In the context of discrete-choice, the model provides falsifiable restrictions which describe how the CCPs should react to the entry of a new product. It is not necessary that the econometrician knows $\phi$, but knowing $\phi$ means $\alpha$ is (over-)identified.
Abstractly, $\phi$ is defined as the choice probability when no information is acquired. A natural assumption is to take $\phi$ to be 1 divided by the number of characteristics available in the market, but $\phi$ could also be other things. In markets where people who acquire little to no information may be identified, $\phi$ is identified.
(Over-)Identification of $\alpha$ via entry. Assume the market has $J$ characteristics $\xi_{1},...,\xi_{J}$ and $\phi$ is known. Consider entry by a new product which changes $\phi$ to $\phi^{\prime}$, which induces a change in the marginal $P$ to $P^{\prime}$. For any $\theta$, we have that
$\frac{P(\xi|\theta)}{P^{\prime}(\xi|\theta)}=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}}{\phi^{\prime}(\xi)^{\alpha}P^{\prime}(\xi)^{1-\alpha}}\frac{Z^{\prime}(\theta;P^{\prime}|_{\xi})}{Z(\theta;P|_{\xi})}$
For any pair $\xi_{1},\xi_{2}\in X$, we have
$\frac{P(\xi_{1}|\theta)}{P^{\prime}(\xi_{1}|\theta)}\frac{P^{\prime}(\xi_{2}|\theta)}{P(\xi_{2}|\theta)}=\frac{\phi(\xi_{1})^{\alpha}P(\xi_{1})^{1-\alpha}}{\phi^{\prime}(\xi_{1})^{\alpha}P^{\prime}(\xi_{1})^{1-\alpha}}\frac{\phi^{\prime}(\xi_{2})^{\alpha}P^{\prime}(\xi_{2})^{1-\alpha}}{\phi(\xi_{2})^{\alpha}P(\xi_{2})^{1-\alpha}}$
That is, an over-identifying feature of the model is that we can construct a ratio that depends on $\theta$ on the left-hand side but not on the right-hand side. We can test whether it is true that for all pairs of $\xi$, the left-hand side is constant w.r.t. $\theta$. If this is not true, the model is easily rejected. What’s more, testing this restriction does not require any knowledge of $\phi$, which does not appear on the left-hand side. However, knowing $\phi$ naturally pins down $\alpha$.
First-Step Orthogonality and Its Implications
First-step orthogonality is a property of moments of the form $\int f_{P}(\omega)\>dP(\omega)$, where the integrand $f$ depends on the probability with respect to which we are integrating. Loosely, it states that a perturbation in $P$ locally affects the moment only through the re-weighting $dP$. It constitutes a sufficient condition for the Gibbs property and additive separability. It is a relatively unique property of entropy-based costs which essentially comes from the fact that $\frac{d}{dx}\log x=1/x$.
Notation. Let $\cal P$ be the set of Bayes-plausible feasible probabilities. Let $H\in{\cal P}$. Define a path $P_{\varepsilon}:=(1-\varepsilon)P+\varepsilon H$. Clearly, $P_{\varepsilon}$ is defined on $\varepsilon\in[l,r]$ for some $l\leq 0$ and $r\geq 1$, and $P_{\varepsilon}\subset{\cal P}$.
Lemma 6 (First-step orthogonality).
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}E_{P_{t}}[Y(\xi,\theta;P_{\varepsilon})]=0$ for $t\in[l,r)$
Theorem 7 (First-step orthogonality II).
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}U(P_{\varepsilon})=E_{H}[Y(\xi,\theta;P_{t})]-E_{P}[Y(\xi,\theta;P_{t})]$ for $t\in[l,r)$
Corollary 7.1 (Gibbs property).
If $P$ maximizes $U$ then $\xi\mapsto Y(\xi,\theta;P)$ is constant a.s.
Proof.
We have $\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}U(P_{\varepsilon})=E_{\theta\sim\mu}\left[\int Y(\xi,\theta;P)\>d(H-P)(\xi|\theta)\right]$. For every $\theta$ on which the property is not satisfied, we can take $H(\xi|\theta)$ to place mass only on the points where $Y$ is largest such that $\int Y(\xi,\theta;P)\>d(H-P)(\xi|\theta)>0$. ∎
By the same token if $P$ maximizes $U$ subject to a marginal constraint $P|_{\xi}=\nu$, then $Y$ should be additively separable. More on this later.
Lemma 8.
Let $\alpha>0$. If $P$ maximizes $U$ then the density $\frac{dP(\xi)}{d\phi(\xi)}$ does not get arbitrarily close to 0 and does not become arbitrarily large.
Schrödinger Potentials
The problem faced by the DM can be separated into two problems: i) maximizing the objective eq.˜1 subject to a further marginal constraint $P|_{\xi}=\nu$ and ii) solving for the optimal marginal $\nu$. We observe here that the former is an instance of the entropic optimal transport problem. The constrained optimal marginal is called the Schrödinger bridge from $\nu$ to $\mu$.
In lemma˜8, we established that if $P$ maximizes $U$ subject to feasibility and Bayes plausibility, then $P|_{\xi}\sim\phi$ and $dP(\xi)/d\phi(\xi)$ is upper bounded and lower bounded away from zero. So in this section, we assume $P|_{\xi}$ is restricted to a marginal $\nu$ for which $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from zero.
Define ${\cal Q}(\nu)=\{P\in\Delta(X\times\Theta):P|_{\xi}=\nu,P|_{\theta}=\mu\}$. We let
$V(\nu)=\sup_{P\in{\cal Q}(\nu)}U(P)$
denote the value of imposing the constraint $P|_{\xi}=\nu$.
We can rewrite U(P)-U(P) as:
$-U(P)$ $=\int-{u(\xi,\theta)}+\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)\>dP+\int\log\left(\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
and turn to minimizing
$-V(\nu)=\inf_{P\in{\cal Q}(\nu)}\int c(\xi,\theta)\>dP-D_{KL}(P\|\nu\otimes\mu)$ (4)
where $c(\xi,\theta)=-{u(\xi,\theta)}+\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)\in L^{\infty}(\nu\otimes\mu)$.
This is called the Entropic Optimal Transport (EOT) problem. [^3] For simplicity, we may assume $c(\xi,\theta)$ is totally bounded, since the worse case scenario choice of $\nu$ can be set to $\phi$, in which case c=uc=-u, which we assumed to be totally bounded. The following theorems concern properties of EOT and can be found in [6], up to notational changes.
Theorem 9 (Existence and uniqueness of potentials).
If $P$ solves eq.˜4, then there exist functions $a_{\nu}\in L^{1}(\nu),-b_{\nu}\in L^{1}(\mu)$ such that
$\frac{dP}{d(\nu\otimes\mu)}=e^{-a_{\nu}(\xi)-b_{\nu}(\theta)-c(\xi,\theta)}=e^{u(\xi,\theta)-a_{\nu}(\xi)-b_{\nu}(\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}$ (5)
These are called the Schrödinger potentials. Conversely, if there exists potentials such that eq.˜5 holds, then $P$ must be optimal. Moreover, potentials are unique up to translation. [^4] That is, for any two pairs of potentials $(a_{\nu},b_{\nu})$ and $(a_{\nu}^{\prime},b_{\nu}^{\prime})$, we have $a_{\nu}-a^{\prime}_{\nu}=b_{\nu}^{\prime}-b_{\nu}$.
Corollary 9.1.
If $P$ solves eq.˜4 then $Y(\xi,\theta;P)=a_{\nu}(\xi)+b_{\nu}(\theta)$ and so $U(P)=E_{\nu}[a_{\nu}(\xi)]+E_{\mu}[b_{\nu}(\theta)]$
Theorem 10 (Schrödinger Equations).
$a_{\nu}$ and $b_{\nu}$ constitute a pair of Schrödinger potentials if and only if they satisfy
$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{e^{b_{\nu}(\theta)}}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}}\>d\nu(\xi)$
Ergo,
$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}}\>d\nu(\xi)}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{b_{\nu}(\theta)}}\>d\mu(\theta)}\>d\nu(\xi)$
Theorem 11 (Duality).
$V(\nu):=\sup_{P:\ P|_{\xi}=\nu,P|_{\theta}=\mu}U(P)$
$V(\nu)$ $=\inf_{a\in L^{1}(\nu),\ b\in L^{1}(\mu)}\int a\>d\nu+\int b\>d\mu+\iint e^{-a-b-c}\>d\nu\>d\mu-1$
$=\inf_{a\in L^{1}(\nu),\ b\in L^{1}(\mu)}\int a\>d\nu+\int b\>d\mu+\iint e^{u(\xi,\theta)-a-b}\>d\phi^{\alpha}d\nu^{1-\alpha}\>d\mu-1$
Note that because $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from 0, by Hölder’s equality with coefficients $p=1$ and $q=\infty$, we can interchange $a\in L^{1}(\nu)$ with $a\in L^{1}(\phi)$.
Corollary 11.1.
Consider a path $\nu_{\varepsilon}$ such that at $\varepsilon=0$, $\varepsilon\mapsto\frac{d\nu_{\varepsilon}(\xi)}{d\phi(\xi)}$ is continuous for $\phi$-a.e. $\xi$, and $d\nu_{\varepsilon}/d\phi$ is uniformly bounded away from 0 and $\infty$ on $\varepsilon\in(-r,r)$. Then $\varepsilon\mapsto V(\nu_{\varepsilon})$ is continuous at 0.
In the discrete case, the Schrödinger potentials can be viewed as the Lagrange multipliers for the marginal constraints. In other words, $a_{\nu}(x)$ can be viewed as the rate of increase in $V$ when adding an infinitessimal mass of $x$ to $\nu$. This is because
$V(\nu)=\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
so in the discrete case, applying the Envelope theorem and normalizing $\int a_{\nu}\>d\nu=0$ gives
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}V\bigg{(}(1-\varepsilon)\nu+\varepsilon\delta_{x}\bigg{)}=a_{\nu}(x)$
The continuous case is more involved and discussion is deferred to the appendix.
The Jensen Upper Bound of the Marginal Value
If one writes out the objective – a function of the joint $P$ – swaps the log with the first integral, one obtains a function of the marginal $P|_{\xi}$. When $\alpha\in(0,1)$ this function is strictly convex, and, by Jensen’s Inequality, upper bounds $V$. We show that their maxima coincide, meaning that solving for the optimal marginal is a relatively simple convex optimization problem.
Define $f(\nu)=E_{\theta\sim\mu}\left[\log\left(E_{\xi\sim\phi}\left[e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}\right]\right)\right]\equiv E_{\theta\sim\mu}[\log(Z(\theta;\nu)]$
Theorem 12.
$f(P|_{\xi})\geq U(P)$
Theorem 13.
$f$ is strictly concave when $\alpha\in(0,1)$ and therefore has a unique maximum.
Theorem 14.
If $P$ maximizes $U$ then $f(P|_{\xi})=U(P)$.
Proof.
This follows from plugging in the weighted MNL formula. ∎
Theorem 15.
The maximum of $f$ and $U$ coincide. That is, if $\nu$ maximizes $f$, then there exists a $P$ with $P|_{\xi}=\nu$ which maximizes $U$.
Corollary 15.1.
The maximum of $U$ is unique.
Conclusion
We study a model of rational inattention to states $\theta$ and hedonic characteristics $\xi$. The DM pays the standard linear-in-Shannon cost for the mutual information contained in the joint distribution over $(\xi,\theta)$. In addition, some choices are intrinsically more common, as represented by the prior $\phi$, and so the DM pays a cost for diverging from $\phi$.
The DM’s conditional choice probability satisfies a weighted multinomial logit:
$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(\xi,\theta)}}{Z(\theta;P)}$
In the state-action model, the DM has full information about their actions and is only learning about the state, and so narrows down the set of choices to a “consideration set.” Because there are many possible optimal choices for marginals $P(a)$, there may not be a unique solution. By contrast, in our model, the DM does not observe the hedonic characteristics of their selections; because it gets increasingly costly to rule out characteristics with increasing certainty, the DM never rules out any choice. Thus, the marginal lies in the interior and, because of strict concavity, is unique.
Appendix
Omitted proofs
Proof of lemma˜1.
Let $P(\cdot|\theta)$ be any disintegration kernel. To see that $P(\cdot|\theta)\ll\phi$, $\mu$-a.s., let $\phi(A)=0$. Then,
$P|_{\xi}(A)=\int P(A|\theta)\>d\mu(\theta)=0$
which implies that $P(A|\theta)=0$ $\theta$-a.s. Therefore, $dP(\xi|\theta)/d\phi(\xi)<\infty$ $\phi\otimes\mu$-a.e. To see $P\ll\phi\otimes\mu$, let $(\phi\otimes\mu)(S)=0$. Then,
$P(S)=\int\textbf{1}_{S}\>dP=\iint\textbf{1}_{S}\frac{dP(\xi|\theta)}{d\phi(\xi)}\>d\phi(\xi)\>d\mu(\theta)$
$\textbf{1}_{S}=0$ $\phi\otimes\mu$-a.e., so $P(S)=0$. So, $dP/d(\phi\otimes\mu)<\infty$, $\phi\otimes\mu$-a.e. ∎
Proof of lemma˜2.
Since $P\ll\phi\otimes\mu$, by the Radon-Nikodym theorem, there is an equivalence class of a.e.-equal densities which are a.e.-finite. By construction, no matter which element of the equivalence class one takes,
$\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\mu(\theta)<\infty\text{; $\phi$-a.s.}$ $\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\phi(\xi)=1\text{; $\mu$-a.s.}$ $\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}<\infty\text{; $\phi\otimes\mu$-a.s.}$
so on the null set of points and fibers where these almost-sure properties do not hold, we can re-define $\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}$ to 1, so that the properties now hold everywhere, not just a.e. Since we are changing the function on a null set, it is still in the equivalence class. Then, we define the other densities as follows:
$\frac{dP(\xi|\theta)}{d\phi(\xi)}=\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}$ $\frac{dP(\xi)}{d\phi(\xi)}=\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\mu(\theta)$ $\frac{dP(\theta)}{d\mu(\theta)}=\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\phi(\xi)$
and
$\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}=\begin{cases}\frac{\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}}{\int\frac{dP(\xi,t)}{d(\phi\otimes\mu)(\xi,t)}\>d\mu(t)}&\text{if the denominator is positive}\\ 1&\text{otherwise}\end{cases}$
It can then be seen that the kernels defined $P(A|\theta):=\int_{A}\frac{dP(\xi|\theta)}{d\phi(\xi)}\>d\phi(\xi)$ and $P(B|\xi)=\int_{B}\frac{dP(\theta|\xi)}{d\mu(\theta)}\>d\mu(\theta)$ are valid disintegration kernels. ∎
Proof of lemma˜6.
We have
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}\frac{dP_{t}(\theta|\xi)}{d\mu(\theta)}\log\left(\frac{dP_{\varepsilon}(\theta|\xi)}{d\mu(\theta)}\right)=\frac{d(H-P)(\theta|\xi)}{d\mu(\theta)}\in L^{1}(\mu)$
for all $t\in[0,1]$. The integral of which is
$\iint\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}\frac{dP_{t}(\theta|\xi)}{d\mu(\theta)}\log\left(\frac{dP_{\varepsilon}(\theta|\xi)}{d\mu(\theta)}\right)\>d\mu(\theta)\>dP_{t}(\xi)=\iint\frac{d(H-P)(\theta|\xi)}{d\mu(\theta)}\>d\mu(\theta)\>dP_{t}(\xi)=0$
From a similar calculation, one obtains
$\int\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}\frac{dP_{\varepsilon}(\xi)}{d\phi(\xi)}\log\left(\frac{dP_{\varepsilon}(\xi)}{d\phi(\xi)}\right)\>d\phi(\xi)=0$
The exchangeability of the integral and the derivative follows from [3, Theorem 2.27]; in particular, the sufficient condition is that the absolute value of the derivative of the integrand is bounded by an element of $L^{1}$. The result follows by construction of $Y$. ∎
Proof of theorem˜7.
By the product rule,
$\displaystyle\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}U(P_{\varepsilon})$ $=\lim_{\varepsilon\downarrow t}\frac{1}{\varepsilon-t}\left[\int Y(\xi,\theta;P_{\varepsilon})\>dP_{\varepsilon}-\int Y(\xi,\theta;P_{t})\>dP_{t}\right]$
$=\lim_{\varepsilon\downarrow t}\frac{1}{\varepsilon-t}\left[\int Y(\xi,\theta;P_{\varepsilon})\>dP_{\varepsilon}-\int Y(\xi,\theta;P_{\varepsilon})\>dP_{t}+\int Y(\xi,\theta;P_{\varepsilon})\>dP_{t}-\int Y(\xi,\theta;P_{t})\>dP_{t}\right]$
$=\left(\lim_{\varepsilon\downarrow t}\frac{1}{\varepsilon-t}\left[\int Y(\xi,\theta;P_{\varepsilon})\>dP_{\varepsilon}-\int Y(\xi,\theta;P_{\varepsilon})\>dP_{t}\right]\right)+\underbrace{\displaystyle\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}E_{P_{t}}\left[Y(\xi,\theta;P_{\varepsilon})\right]}_{=0}$
$=\lim_{\varepsilon\downarrow t}\int Y(\xi,\theta;P_{\varepsilon})\>d(H-P)=\int Y(\xi,\theta;P_{t})\>d(H-P)$
Proof of lemma˜8.
Let $\underline{u}\leq u(\xi,\theta)\leq\overline{u}$. For contradiction, suppose for every $M\in\mathbb{R}$, there was some set $A_{M}$ of mass $\phi(A_{M})>0$ where $\alpha\log\left(dP(\xi)/d\phi(\xi)\right)>M$. Then, since $Y(\xi,\theta;P)=b(\theta)$, $P$-a.s., then it must be the case that for $P(\cdot|\theta)$-a.e. $\xi\in A_{M}$,
$b(\theta)=u(\xi,\theta)-\alpha\log\left(\frac{dP(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)>M+u(\xi,\theta)-\log\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)$
meaning for $P$-a.e. $(\xi,\theta)$,
$\frac{dP(\xi|\theta)}{dP(\xi)}>e^{M+u(\xi,\theta)-b(\theta)}$
which implies
$1=\int\frac{dP(\xi|\theta)}{dP(\xi)}\>d\mu(\theta)>e^{M}\int e^{u(\xi,\theta)-b(\theta)}\>d\mu(\theta)$
for every $M$. This implies that $\int e^{-b(\theta)}\;d\mu(\theta)=0$ thus raising the contradiction.
Along the same lines, suppose for every $M\in\mathbb{R}$, there was some set on which $\alpha\log\left(dP(\xi)/d\phi(\xi)\right)<-M$. We have $b(\theta)<-M+u(\xi,\theta)-\log\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)$, and thus
$1=\int\frac{dP(\xi|\theta)}{dP(\xi)}\>dP(\xi)<e^{-M}e^{b(\theta)}\int e^{u(\xi,\theta)}\>dP(\xi)\leq e^{-M}e^{b(\theta)}e^{\overline{u}}$
so $e^{b(\theta)}=\infty$ everywhere, which also raises a contradiction. ∎
Proof of corollary˜11.1.
Suppose $\gamma,\nu$ are two marginals, and without loss $V(\gamma)\geq V(\nu)$. We can upper bound $V(\gamma)$ by using the “wrong” Schrödinger potentials in the dual objective. In particular, let $\gamma=\nu_{\varepsilon}$
$|V(\nu_{\varepsilon})-V(\nu)|\leq\bigg{|}\int a_{\nu}\>d(\nu_{\varepsilon}-\nu)+\iint e^{u(\xi,\theta)-a_{\nu}-b_{\nu}}\left[\left(\frac{d\nu_{\varepsilon}}{d\phi}\right)^{1-\alpha}-\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\right]\>d\phi\>d\mu-1\bigg{|}$
But this is continuous w.r.t. $\varepsilon$. Since $\frac{d\nu_{\varepsilon}}{d\phi}$ is uniformly bounded away from 0 and $\infty$ in a neighbourhood of $\varepsilon=0$, we apply the Dominated Convergence Theorem to get $V(\nu_{\varepsilon})\to V(\nu)$. ∎
Proof of theorem˜12.
By Jensen’s inequality,
$U(P)$ $=E_{\theta\sim\mu}\left[\int\log\left(e^{u(\xi,\theta)}\left(\frac{dP(\xi)}{d\phi(\xi)}\right)^{-\alpha}\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)^{-1}\right)dP(\xi|\theta)\right]$
$\leq E_{\theta\sim\mu}\left[\log\left(\int e^{u(\xi,\theta)}\left(\frac{dP(\xi)}{d\phi(\xi)}\right)^{-\alpha}\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)^{-1}dP(\xi|\theta)\right)\right]$ (Jensen’s Inequality)
$=E_{\theta\sim\mu}\left[\log\left(\int e^{u(\xi,\theta)}\left(\frac{dP(\xi)}{d\phi(\xi)}\right)^{1-\alpha}d\phi(\xi)\right)\right]$ (change of measure)
$=f(P|_{\xi})$
Proof of theorem˜13.
Let $\gamma,\nu$ be feasible probability measures and $\beta\in(0,1)$. Since $\alpha\in(0,1)$, $x\mapsto x^{1-\alpha}$ is strictly concave, it follows
$Z(\theta;\beta\gamma+(1-\beta)\nu)$ $=\int\left(\beta\frac{d\gamma(\xi)}{d\phi(\xi)}+(1-\beta)\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}e^{u(\xi,\theta)/\lambda}d\phi(\xi)$
$>\beta\int\left(\frac{d\gamma(\xi)}{d\phi(\xi)}\right)^{1-\alpha}e^{u(\xi,\theta)/\lambda}d\phi(\xi)+(1-\beta)\int\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}e^{u(\xi,\theta)/\lambda}d\phi(\xi)$
$=\beta Z(\theta;\gamma)+(1-\beta)Z(\theta;\nu)$
Moreover, $\log$ is strictly concave and strictly increasing, so
$\log\left(Z(\theta;\beta\gamma+(1-\beta)\nu)\right)>\log\left(\beta Z(\theta;\gamma)+(1-\beta)Z(\theta;\nu)\right)>\beta\log(Z(\theta;\gamma))+(1-\beta)\log(Z(\theta,\nu))$
From which we conclude that
$f\left(\beta\gamma+(1-\beta)\nu\right)$ $=\int\log\left(Z(\theta;\beta\gamma+(1-\beta)\nu)\right)\>d\mu(\theta)$
$>\beta\int\log(Z(\theta;\gamma))\>d\mu(\theta)+(1-\beta)\int\log(Z(\theta;\nu))\>d\mu(\theta)$
$=\beta f\left(\gamma\right)+(1-\beta)f\left(\nu\right)$
as needed. ∎
Proof of theorem˜14.
If $\nu$ maximizes $f$, then a first-order condition is that for any $\gamma$ which defines a path $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$, we have $df(\nu_{\varepsilon})/d\varepsilon\leq 0$ at 0, which implies:
$0$ $\geq\frac{df(\nu_{\varepsilon})}{d\varepsilon}\bigg{|}_{\varepsilon=0}\geq\int\frac{\int(1-\alpha)e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\frac{d(\gamma-\nu)(\xi)}{d\phi(\xi)}d\phi(\xi)}{\int e^{u(\xi,\theta)}d\phi^{\alpha}d\nu^{1-\alpha}}d\mu=\int\frac{\int e^{u(x,\theta)}\left(\frac{d\nu(x)}{d\phi(x)}\right)^{-\alpha}\>d\gamma(x)}{\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\>d\nu(\xi)}-1\>d\mu$
Since this holds for all $\gamma\ll\phi$, it follows for $\phi$-a.e. $x$,
$\int\frac{e^{u(x,\theta)}\left(\frac{d\nu(x)}{d\phi(x)}\right)^{-\alpha}}{\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\>d\nu(\xi)}-1\>d\mu\leq 0$
(To show this, assume for contradiction that there is a set $A$ with $\phi(A)>0$ such that the converse is true. Then, take $\gamma$ such that $d\gamma(x)/d\phi(x)>0$ iff $x\in A$. Integrating the left-hand side of the above inequality w.r.t. $\gamma$ leads to the contradiction.) Since the left-hand side must sum to unity when integrating w.r.t. $\nu$, it further follows that
$\int\frac{e^{u(x,\theta)}\left(\frac{d\nu(x)}{d\phi(x)}\right)^{-\alpha}}{\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\>d\nu(\xi)}-1\>d\mu=0$
for a.e. $x$. Then, $a(\xi)=0$ and $b(\theta)=\log\left(\int e^{u(\xi,\theta)}\>d\phi(\xi)^{\alpha}\>d\nu(\xi)^{1-\alpha}\right)$ satisfy the Schrödinger equations for the marginal $\nu$. Thus, it follows if $\nu$ maximizes $f$,
$dP(\xi,\theta)=\frac{e^{u(\xi,\theta)}d\phi(\xi)^{\alpha}d\nu(\xi)^{1-\alpha}}{e^{b(\theta)}}$
is not only a valid probability measure, but also $P|_{\xi}=\nu$, $P|_{\theta}=\mu$, and $P$ maximizes $U$ subject to those constraints. ∎
Proof of corollary˜15.1.
If $P,P^{\prime}$ both maximize $U$, then $P|_{\xi},P^{\prime}|_{\xi}$ both maximize $f$. Since $f$ is strictly concave, it follows that the marginals are equal: $P|_{\xi}=P^{\prime}|_{\xi}=\nu$. Because the joint must satisfy the MNL formula, it follows $P$ and $P^{\prime}$ are just versions of one another. ∎
The Envelope
As suggested, the Envelope Theorem provides a nice heuristic interpretation for $a_{\nu}(x)$. It is left to show, however, that $V(\nu_{\varepsilon})$ or $V(\nu_{\varepsilon})$ can be written as the supremum of a function whose derivative is equicontinuous [5].
If $b_{\nu}$ is a Schrödinger potential, then plugging in $a_{\nu}=\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b_{\nu}}\>d\mu\right)$ gives
$V(\nu)=\int\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b_{\nu}}\>d\mu\right)\ d\nu-\int b_{\nu}\>d\mu$
Normalize $u$ such that $\underline{u}\geq 0$. We are allowed to perform one normalization on $b$, so assume $\int b_{\nu}\>d\mu=0$. Then,
$V(\nu)=-\alpha D_{KL}(\nu\|\phi)+\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)\>d\nu$
By construction of $V(\nu)=\sup_{P\in{\cal Q}(\nu)}U(P)$,
$\overline{u}-\alpha D_{KL}(\nu\|\phi)\geq V(\nu)$
so that
$\overline{u}\geq\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)\>d\nu$
Because $\underline{u}>0$, we have
$\log\left(\int e^{-b_{\nu}}\>d\mu\right)\leq\log\left(\int e^{\underline{u}-b_{\nu}}d\mu\right)\leq\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)d\nu\leq\overline{u}$
Define $B$ as
$B=\bigg{\{}b\in L^{1}(\mu):\int b\>d\mu=0,\ \log\left(\int e^{-b}\;d\mu\right)\leq\overline{u}\bigg{\}}$
The dual can be written as
$-V(\nu)$ $=\sup_{a\in L^{1}(\phi),\ b\in L^{1}(\mu)}\int-a\>d\mu+\int-b\>d\nu-\iint e^{u(\xi,\theta)-a-b}\>d\phi^{\alpha}d\nu^{1-\alpha}\>d\mu+1$
$=\sup_{b\in B}\int-\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b}\>d\mu\right)\ d\nu$
$=\alpha D_{KL}(\nu\|\phi)+\sup_{b\in B}\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu$
since $B$ imposes no binding restrictions.
Define $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$ such that $d\nu_{\varepsilon}/d\phi$ is bounded uniformly away from 0 and $\infty$ on $\varepsilon\in(-r,r)$. Define $g:B\times(-r,r)\to\mathbb{R}$ by
$g(b,\varepsilon)=\alpha D_{KL}(\nu\|\phi)+\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu$
Then,
$-V(\nu_{\varepsilon})=\sup_{b\in B}g(b,\varepsilon)$
By first-step orthogonality, $\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}D_{KL}(\nu_{\varepsilon}\|\phi)=\int\log\left(\frac{d\nu_{t}}{d\phi}\right)\>d(\gamma-\nu)$, and so
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)=\int\alpha\log\left(\frac{d\nu_{t}}{d\phi}\right)-\log\left(\int e^{u-b}\>d\mu\right)\ d(\gamma-\nu)$
By assumption, $\log(d\nu_{t}/d\phi)$ is uniformly bounded. By Jensen’s inequality,
$\log\left(\int e^{u-b}\>d\mu\right)\geq\int u-b\>d\mu=\underline{u}$
Conversely, since $\int e^{-b}\>d\mu\leq e^{\overline{u}}$, by Holder’s inequality
$\int e^{u-b}\>d\mu\leq e^{\overline{u}}\int e^{-b}\>d\mu=e^{2\overline{u}}$
so $\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)$ is uniformly bounded across $b\in B$.
Furthermore, notice that for any $b$,
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)-\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=s}g(b,\varepsilon)=\int\alpha\log\left(\frac{d\nu_{t}}{d\nu_{s}}\right)\>d(\gamma-\nu)$
i.e. it does not depend on $b$. Thus, $\bigg{\{}\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)\bigg{\}}_{b\in B}$ is equicontinuous.
Lastly,
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}g(b_{\nu},\varepsilon)=\int-\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b_{\nu}}\>d\mu\right)\ d(\gamma-\nu)=\int-a_{\nu}\>d(\gamma-\nu)$
Thus,
$\frac{dV(\nu_{\varepsilon})}{d\varepsilon}\bigg{|}_{\varepsilon=0}=\int a_{\nu}\>d(\gamma-\nu)$
If $X$ is a metric space with full support and $a_{\nu}$ is continuous and bounded, then we can construct a sequence of probabilities $\gamma_{n}$ which converges weakly to the Dirac measure $\delta_{x}$, so that
$\frac{dV(\nu_{\varepsilon,n})}{d\varepsilon}\bigg{|}_{\varepsilon=0}\to_{n}\int a_{\nu}\>d(\delta_{x}-\nu)=a_{\nu}(x)-\int a_{\nu}\>d\nu$
Sinkhorn’s algorithm
In the discrete case, the problem of finding Schrödinger potentials such that eq.˜5 holds is equivalent to the matrix scaling problem. Formally, we can re-write eq.˜5 in the discrete case as
$P(\xi,\theta)=e^{-a_{\nu}(\xi)}\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]e^{-b_{\nu}(\theta)}$
If we think of the middle term as a $|X|\times|\Theta|$ matrix $A$, the Schrödinger potentials as diagonal matrices, and $P$ as a $|X|\times|\Theta|$ matrix such that the marginals are $\nu$ and $\mu$ respectively, we can re-formulate the problem as follows: given a matrix A and a pair of probability vectors $\boldsymbol{\nu}$, $\boldsymbol{\mu}$, find diagonal matrices $\textbf{D}_{1},\textbf{D}_{2}$ such that for $\textbf{P}:=\textbf{D}_{1}\textbf{A}\textbf{D}_{2}$ is a probability matrix whose columns sum to $\sum_{\xi}P_{\xi\theta}=\mu_{\theta}$ and rows sum to $\sum_{\theta}P_{\xi\theta}=\nu_{\xi}$.
Theorem 16 (Sinkhorn’s algorithm).
Set $a_{t}=0$. For $t\geq 0$, set
$b_{t}(\theta)=\log\left(\int e^{u(\xi,\theta)-a_{t}(\xi)}\>d\phi^{\alpha}\>d\nu^{1-\alpha}\right)$
$a_{t+1}(\theta)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)$
We may define $P_{2t}$ by via the potentials $a_{t}$, $b_{t}$, and $P_{2t-1}$ via the potentials $a_{t},b_{t-1}$.
Then:
  1. 1.
    $a_{t}$ and $b_{t}$ converge pointwise and in their respective $L^{p}$-norms, $1\leq p<\infty$
  2. 2.
    $D_{KL}(P\|P_{t})\to 0$
  3. 3.
    $D_{KL}(P_{t}\|R)\to D_{KL}(P\|R)$
  4. 4.
    $P_{t}\to P$ in total variation
Proof.
By assumption, $u$ and $\log\frac{d\nu}{d\phi}$ are totally bounded. This then follows from [6, Theorem 6.15]. ∎

### v2

@@
 Abstract
-We study a generalization of the state-action rational inattention model with two dimensions of uncertainty: states and hedonic characteristics.
-The resulting conditional choice probability takes a familiar weighted multinomial logit form, but in contrast to the state-action model, is unique and is not a corner solution.
-The model imposes testable restrictions on the behavior of choice probabilities in markets with product entry.
-The optimal information policy can be found by optimizing subject to a marginal constraint, then finding the optimal marginal.
-The solution to the former is the Schrödinger bridge of an entropic optimal transport problem.
-The latter is solved by maximizing a strictly concave function (over a convex set) which envelops the objective.
+We introduce a rational inattention model which produces a unique, interior, weighted multinomial logit conditional choice probability for an agent who acquires costly information about the hedonic characteristics (e.g.
+whether an insurance contract has high coverage) of their choices and about their payoff-relevant states (e.g.
+their risk of incurring a loss).
+As usual, the objective is to choose a joint distribution subject to one marginal constraint (“Bayes plausibility”).
+We approach the problem by re-writing it in terms of an inner problem of maximizing over two constraints and an outer problem of choosing the “optimal constraint.” The inner problem is a Schrödinger bridge problem.
+The outer problem is strictly concave.
 Introduction
-Consumer choice typically contains two dimensions of uncertainty: one concerns the product, and the second concerns utility-relevant future states of the world.
-When buying insurance, for example, the consumer faces uncertainty surrounding the contents of each insurance contract and their own future risk.
-An increasingly popular framework for studying decision-making under uncertainty is the state-action rational inattention model [4].
-In it, the decision-maker (DM) commits to an information policy consisting of a joint distribution $P$ over actions and states, subject to a Bayes plausibility constraint.
-The DM seeks to maximize expected utility less an information cost which is linear in the Shannon information between state and action.
-A relationship between the optimal conditional choice probability and the optimal marginal probability is known; however, the optimal marginal is not necessarily unique, and therefore neither is the optimal conditional.
-We present a natural extension of the state-action model, which we call the state-characteristic model.
-In this model, the DM faces an exogenous prior distribution $\phi$ over characteristics $\xi$ – reflecting the inherent ubiquity of certain characteristics – and commits to a joint distribution $P$ over characteristics $\xi$ and states $\theta$.
-Mathematically, our model differs in only one respect: in addition to the usual Shannon information cost, we add a cost of diverging from $\phi$, interpretable as information gain from $P$ over $\phi$.
-This models, for instance, the fact that in a market full of lemons, one would need to acquire information to ensure that the likelihood of buying lemons is low.
-The dual information costs account for the two dimensions of information gain that is observed by the econometrician when they observe the joint distribution over $(\xi,\theta)$.
-First, the econometrician observes there is information gain from $P(\xi|\theta)$ over $\phi(\xi)$ – that is, the choice distribution over characteristics is not what it would be if consumers chose at random.
-Second, the econometrician observes that the consumer’s choice of coverage level provides information about what their future losses will be – which differs from the prior $\mu$.
-In section˜2, we motivate the need for the state-characteristic model by showing that the state-action model cannot account for both dimensions of information observed by the econometrician: modeling characteristics as actions fails to account for information gain from $P(\xi|\theta)$ over $\phi(\xi)$, while modeling characteristics as stochastic fails to account for information gain from $P(\theta|\xi)$ over $\mu(\theta)$.
-The state-characteristic model can be thought of as an interpolation between these two approaches which includes both information costs.
-We present the model in full generality in section˜3.
-In our model, the relative costs for each dimension of information is parametrized by $\alpha\in[0,1]$.
-Our model thus nests the two aforementioned versions of the state-action model discussed in section˜2 at $\alpha=0$ and $\alpha=1$.
-As results have already been established for these cases, we focus on the case $\alpha\in(0,1)$.
-Our main result, presented in section˜4, is very similar to the famous weighted multinomial logit (MNL) result [4]. Whereas their optimal conditional choice probability (CCP) satisfied
-$P(a|\theta)=\frac{P(a)e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
-for some normalizing partition function $Z$, our CCP satisfies
-$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
-One major advantage that our model ($\alpha\in(0,1)$) has is it always admits a unique solution whereas state-action ($\alpha=0$) does not.
-Therefore, even in cases where state-action seems to better model the situation, the state-characteristic model may nevertheless be useful for selecting from multiple solutions.
-Another distinction is that in state-action model, the DM forms (possibly non-unique) optimal “consideration sets” [1] – such that actions outside the set have $P(a)=0$ – whereas a key feature of the state-characteristic model is that no characteristic is completely ruled out.
-This reflects the idea that consumers typically content themselves with being pretty sure of certain things; being truly 100% sure of anything requires an irrational amount of effort.
-For example, to acquire information as to how sausages are made, the conscious consumer may read the nutrition facts, the concerned consumer might read the ingredients, but nobody burglarizes a sausage factory with a spectrometer to make sure that the sausages are not produced in an environment with trace amounts of soy.
-In lay terms, what the MNL formula states is that the CCP depends on three things.
-The first is the utility.
-The second is the inherent commonality of characteristics $\phi$: the more lemons there are, the more likely it is that you will buy one.
-The third is the marginal probability $P$.
-If you prefer $\xi$ to $\xi^{\prime}$ in 99% of states, then even in the 1% of states where you slightly prefer $\xi^{\prime}$ to $\xi$, you may still end up picking $\xi$ unless you acquire enough information about the state to be quite sure you are in that 1%.
-How much each of these three things matters depends on the parameters $\alpha,\lambda$.
-If $\alpha\lambda$ is large, that means it is harder to learn about the characteristics $\xi$.
-Thus, $P(\xi|\theta)$ will resemble what would happen if you were to choose randomly, which is represented by the prior $\phi(\xi)$.
-If $(1-\alpha)\lambda$ is large, that means it is harder to learn about states $\theta$.
-Thus, $P(\xi|\theta)$ will not differ much from $P(\xi)$, since separating $P(\xi|\theta)$ from $P(\xi)$ means you learned to differentiate $\theta$ from other states.
-If both are large, then $P(\xi)$ is close to $\phi(\xi)$ and the CCPs are close to both.
-Conversely, as $\lambda\to 0$, the CCP begins to look like that of a fully-informed, utility (soft-)maximizing agent.
-Lastly, section˜5 concerns falsifiable restrictions implied by the results.
-In particular, entry of products into the market which generate variation in $\phi$ should generate changes in CCP ratios that are not state-dependent.
-These predictions do not require knowledge of the prior $\phi$ to be tested.
-However, if the prior is known, the relative cost parameter $\alpha$ is over-identified in the process.
-In the finite discrete case, the results may be established by taking Lagrangians.
-In the remaining sections, we derive the results for the general case, since characteristics may be continuous.
-The results hereafter are mostly mathematical in nature, whose utility lies in that they provide (hopefully) interesting and (hopefully) novel insights into the problem, as well as connections to other well-known problems.
-The proofs in these sections are typically deferred to the appendix.
-Section˜6 explores the first-step orthogonality property of Shannon entropy-based inattention objectives and its consequences.
-In brief, the objective (which is a function of the information policy $P$) can be written as an integral in which $P$ appears twice: once in the integrand and once in the form of $dP$.
-While the product rule suggests that perturbing $P$ ought to affect the objective in two ways, first-step orthogonality establishes that it only locally affects the objective in one: via the re-weighting in the form of $dP$.
-First-step orthogonality implies that both state-action and state-characteristic have a nice property whereby when the objective is written as a single integral, the integrand is constant.
-Section˜7 establishes a connection between rational inattention and entropic optimal transport (EOT).
-It turns out that the problem of maximizing the state-characteristic objective subject to a second marginal constraint $P|_{\xi}=\nu$ (on top of Bayes plausibility, which is itself a marginal constraint $P|_{\theta}=\mu$) can be re-cast as a Schrödinger bridge problem [7].
-The problem comes with a dual which consists of finding Schrödinger potentials $a(\xi),b(\theta)$ which satisfy the associated Schrödinger equations.
-From EOT theory, we know that the value as a function of the marginal constraint $\nu$ is additively separable and equal to the sum of (the expectations of) the potentials.
-The potentials have a straightforward Lagrangian interpretation.
-Loosely speaking, the potential associated with the characteristic $\xi$ corresponds to how much the value would change if we were to change the marginal constraint $\nu$ such that some extra mass were added to $\xi$.
-Perhaps the most interesting practical takeaway here is that the optimal CCP subject to two marginal constraints is also a weighted MNL. [4] had established that
-$\text{$P$ is optimal}\implies P(\xi|\theta)\propto P(\xi)e^{u(\xi,\theta)/\lambda}$
-but this does not necessarily imply that
-$\text{$P$ is optimal subject to }P|_{\xi}=\nu\implies P(\xi|\theta)\propto\nu(\xi)e^{u(\xi,\theta)/\lambda}$
-Indeed, it is not true. However, the Schrödinger equations imply that $P$ maximizes the objective subject to $P|_{\xi}=\nu,P|_{\theta}=\mu$ if and only if there exists $a,b$ such that
-$P(\xi|\theta)=\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}e^{[u(\xi,\theta)-a(\xi)-b(\theta)]/\lambda}$
-Thus, $P(\xi|\theta)$ always follows a weighted MNL.
-Finding such an $a,b$ when $\xi,\theta$ are discrete is a case of the matrix scaling problem.
-Sinkhorn’s algorithm can be used to find the Schrödinger potentials.
-We finish by considering the problem of finding the optimal marginal in section˜8.
-In [1], the authors write out a concave function over the set of marginals which helps them establish that the Kuhn-Tucker conditions are sufficient, as opposed to only necessary.
-This function may be obtained by swapping an expectation with a log; by Jensen’s inequality, it upper bounds the objective.
-It turns out that this function has even nicer properties when $\alpha\in(0,1)$.
-In particular, it is strictly concave over a convex set and coincides with the objective at their mutual optima.
-Thus, rather than maximize a more complicated objective, it suffices to solve this much easier convex optimization problem.
+Many welfare-relevant choice problems feature uncertainty to both payoff-relevant states and menu characteristics.
+Consider insurance.
+A household may choose a contract that later appears ex-post “wrong”—too much or too little coverage.
+This could reflect uncertainty about the state (future medical expenditures), but it could also reflect limited attention to the menu (misunderstanding which plan is high coverage, misperceiving deductibles, or failing to screen out dominated options).
+These two failures have different normative implications: policies that improve forecasting of health shocks are not the same as policies that improve comprehension of contract characteristics.
+This paper aims to build upon the state-action rational inattention model [matvejka2015rational], which has become a leading workhorse for disciplined departures from full information.
+In that framework, the decision-maker commits to a joint distribution over actions and states subject to Bayes plausibility, paying a cost linear in Shannon mutual information.
+The model yields a logit-like conditional choice rule and provides a clean mapping from information costs to stochastic choice.
+However, when the economic object of interest is choice among options with uncertain characteristics, the state–action approach confronts two limitations highlighted by the motivating insurance setting.
+First, it does not separate uncertainty about states from uncertainty about characteristics: treating characteristics as actions assumes they are costlessly observed; treating them as part of the state can render prior private information about other state components observationally irrelevant—a problem in decision problems with multi-dimensional uncertainty that we elaborate on in section˜2.
+Second, the state–action objective need not select a unique optimal marginal distribution over actions, and thus may fail to predict a unique conditional choice probability.
+This paper proposes a generalized rational inattention model that directly targets these issues.
+The overarching idea is to model the econometrician’s observed joint distribution over characteristics and states as the result of a information policy which acquires information both (i) to match characteristics to states and (ii) to deviate from the baseline prevalence of characteristics in the environment.
+Formally, the decision-maker chooses an information policy $P$ over characteristics $\xi\in X$ and states $\theta\in\Theta$, subject to Bayes plausibility $P_{\theta}=\mu$ and feasibility $P_{\xi}\ll\phi$, where $\mu$ is the prior over states and $\phi$ is an exogenous prior over characteristics capturing their ubiquity in the market.
+The objective trades off expected utility against two information costs: the usual mutual information between $\xi$ and $\theta$ (learning about states via choice), and an additional KL divergence penalty that prices deviations of the marginal $P_{\xi}$ from $\phi$ (learning to screen the menu).
+A single parameter $\alpha\in[0,1]$ indexes the relative importance of these two margins of attention.
+The model nests the canonical state–action formulation at $\alpha=0$ and a “characteristics-attention” limit at $\alpha=1$.
+This paper makes several contributions.
+First, it presents a tractable and testable stochastic-choice characterization.
+For intermediate $\alpha\in(0,1)$, optimal conditional choice probabilities satisfy a weighted multinomial logit:
+$P(\xi|\theta)\propto\phi(\xi)^{\alpha}\,P_{\xi}(\xi)^{1-\alpha}\,e^{u(\xi,\theta)/\lambda}$ (1)
+with an appropriate normalizing partition function.
+Equation (1) clarifies the economic forces behind stochastic choice: utility tilts choices toward high-payoff characteristics; $\phi$ captures the pull of prevalent characteristics (“lemons are harder to avoid when they are everywhere”); and the endogenous marginal $P_{\xi}$ captures the inertia induced by costly state learning (choices that are optimal in most states become focal unless attention is expended to condition sharply on $\theta$).
+The result parallels the celebrated logit characterization in [matvejka2015rational] while adding a new, economically meaningful weighting by $\phi$.
+Second, it presents a rational inattention model that yields a point prediction.
+In contrast to the state-action model, a major advantage of this model for $\alpha\in(0,1)$ is that it delivers a unique optimal marginal $P_{\xi}$ and hence a unique conditional choice rule.
+Moreover, the solution is interior relative to $\phi$.
+This feature both improves empirical discipline by providing point predictions and rationalizes phenomena that are difficult to accommodate when characteristics are assumed perfectly observed, such as selection of dominated or low-quality options when such options are common.
+Finally, this paper generalizes rational inattention to information policies over non-discrete spaces and presents a novel connection between rational inattention and recent developments in entropic optimal transport.
+Fixing the marginal $P_{\xi}=\nu$ converts the inner problem into an entropic optimal-transport problem (a Schrödinger bridge) between $\nu$ and $\mu$.
+Because the Schrödinger bridge is unique for every $\nu$, the problem simplifies to finding the optimal $\nu$.
+This turns out to be a strictly convex optimization problem.
+Section 2 motivates the model by formalizing why the standard state–action approach cannot simultaneously discipline both observed information gain in $P(\xi|\theta)$ relative to a characteristic baseline and observed informativeness of choices about states.
+Section 3 introduces the state–characteristic model on general Polish spaces, defines the two-part information cost, and discusses the nesting cases $\alpha=0$ and $\alpha=1$.
+Section 4 states the main characterization results: the Gibbs property and the weighted logit formula, along with uniqueness and interiority.
+Section 5 derives testable restrictions and the overidentification logic using entry-driven shifts in $\phi$.
+Sections 68 develop the variational and optimal-transport machinery: first-step orthogonality, the Schrödinger bridge formulation of the constrained problem, and the strictly concave marginal program that delivers existence and uniqueness.
+We conclude with a section 10 which discusses computation in finite spaces.
+Notation.
+We use $X$ to denote the space of characteristics and $\Theta$ to denote the space of states.
+$\xi$ is a generic element of $X$, typically representing the characteristic of a choice of the decision-maker.
+$\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$ is a random vector that represents the characteristics of a menu $A=\{1,...,J\}$.
+$P$ denotes an information policy, typically the optimal one.
+$E_{Q}$ is the expectation with respect to a measure $Q$.
+For a joint distribution $Q$ over random elements $Y$ and $Z$, let $Q_{Y}$ and $Q_{Z}$ denote the respective marginals.
+$Q\ll R$ means that $Q$ is absolutely continuous with respect to $R$.
+$D_{KL}(Q\|R)=E_{Q}\left[\log\left(\frac{dQ}{dR}\right)\right]$ is the information gain of $Q$ over $R$, commonly known as the Kullback-Leibler divergence.
+$I_{Q}(Y,Z)=D_{KL}(Q_{Y,Z}\|Q_{Y}\otimes Q_{Z})$ is the mutual (Shannon) information between random elements $Y\in S$ and $Z\in T$ under joint distribution $Q$.
+$\Delta S$ represents the set of all probability distributions over $S$.
+$\Pi(\cdots)$ is the set of all joint distributions satsifying the marginal restrictions $(\cdots)$.
 Motivation
-State-action presents two options for modeling consumer choice with uncertain characteristics: either treat the characteristics as actions – i.e.
-costlessly observed – or as a state variable, alongside other state variables.
-The latter option, however, presents an identification problem.
-It turns out that a DM who does not observe the other state variables can achieve the same payoff at the same information cost as someone who costlessly observes them.
-Thus, the two behave identically and are observationally equivalent.
-Example.
-Consumers choose between two insurance contracts $j=1,2$.
-They know one is high-coverage $\xi_{j}=H$ and the other is low-coverage $\xi_{j}=L$ but do not know which is which.
-The consumer has also two future health expenditure states, $t\in\{H,L\}$, each with probability $1/2$.
-The consumer receives utility 1 if they choose the ‘correct’ contract (high coverage iff high future expenditures) and 0 otherwise.
-In other words,
-$u(j,\xi_{1},\xi_{2},t)=\textbf{1}\{\xi_{j}=t\}=\begin{cases}1&\xi_{j}=t\\ 0&\xi_{j}\neq t\end{cases}$
-Recall the state-action sequence proceeds as follows: the consumer commits to an information policy – a Bayes plausible joint distribution over actions $j$ and states $(\xi_{1},\xi_{2},t)$.
-Nature then draws an action recommendation and a state.
-Now, suppose consumer 1 observes $t$ before committing to the information policy
-if $\xi_{1}=t$, recommend $j=1$ with probability $p$ and $j=2$ with probability 1p1-p
-if $\xi_{2}=t$, recommend $j=2$ with probability $p$ and $j=1$ with probability 1p1-p
-and consumer 2 does not observe $t$, but commits to the same information policy. Consider:
-  1. 1.
-    Do the two consumers have the same information upon receiving the same recommendation $j$?
-  2. 2.
-    Do the two consumers pay the same information cost?
-  3. 3.
-    Is the expected payoff (averaged over $t$) the same?
-The answer to the first, clearly, is no.
-Consumer 1 knows $t$ and $j$, whereas consumer 2 only gets $j$.
-Consumer 1 knows, with probability $p$, that the coverage level will be $t$, whereas consumer 2 merely knows that with probability $p$, the coverage level will be appropriate for their future health expenditures – but does not know what either are.
-The answer to the second is yes.
-The answer to the third is also yes – for both, their utility will be 1 with probability $p$ and 0 with 1p1-p.
-Ergo, consumer 2 is able to achieve the same expected payoff at the same information cost, despite having less information before committing to the information policy and after receiving the recommendation.
-Worse, for the econometrician, the conditional choice probability for both consumers is the same, so they are indistinguishable.
-What is happening is that both consumers are acquiring the same amount of information about whether $j$ is the “right” choice.
-Consumer 1 already knows $t$ and is doing so by acquiring information about what $j$ is.
-But consumer 2 is doing so without acquiring any information about what $j$ is or why it might be the right choice.
-What the state-action model accounts for is the fact that for each $t$, the probability of choosing $\xi_{j}=t$ differs from $1/2$ – what it would be if no information was acquired; what it doesn’t account for – or rather, treats as costless – is the information about $t$ itself.
-This issue can be formalized as follows.
-Observational equivalence in state-action models when characteristics are state variables.
-In general, could an econometrician determine whether an insurance buyer in the state-action model has a priori private information about their future health expenditures $\theta$?
-Consider the canonical consumer problem setting (see e.g.
-[1]), the action space is $A=\{1,...,J\}$, and the utility vector is $\textbf{u}(\boldsymbol{\xi},\theta)=(u(1,\boldsymbol{\xi},\theta),...,u(J,\boldsymbol{\xi},\theta))$, which is a function of the random variables $j\in A$, $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$, and $\theta\in\Theta$.
-The state-action model predicts that the optimal conditional choice probability is
-$P(j|\boldsymbol{\xi},\theta)=P(j|\textbf{u}(\boldsymbol{\xi},\theta))=\frac{e^{u(j,\boldsymbol{\xi},\theta)}}{\sum_{k\in[J]}e^{u(k,\boldsymbol{\xi},\theta)}}$
-This equation holds regardless of whether $\theta$ is assumed to be known to the consumer a priori or not. Thus, whether $\theta$ is known a priori is not identified in the state-action model.
-Mathematically, we observe that by the law of iterated expectations,
-$I_{P}(j,\textbf{u}):=\int\log\left(\frac{dP(j,\textbf{u})}{dP(j)dP(\textbf{u)}}\right)\>dP=\iint\log\left(\frac{dP(j,\textbf{u}|\theta)}{dP(j|\theta)dP(\textbf{u}|\theta)}\right)\>dP(j,\textbf{u}|\theta)\>dP(\theta)=\int I_{P}(j|_{\theta},\textbf{u}|_{\theta})\>dP(\theta)$
-because $P(j|\theta)=P(j)$. So, the DM maximizes
-$E_{P}[u]-\lambda I_{P}(j,\textbf{u})\equiv E_{P}\bigg{[}E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})\bigg{]}$
-if and only if they maximize $E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})$ for almost every $\theta$.
-In words, the DM solves the state-action rational inattention problem overall if and only if they solve the state-action rational inattention problem in every health state $\theta$.
-One final observation is that $I_{P}(j|_{\theta},\textbf{u}|_{\theta})=D_{KL}(P(\xi|\theta)\|\phi(\xi))$ for a natural prior $\phi$.
-To see this, suppose that out of the $J$ options, $k_{x}$ of them had the characteristic $x\in X$.
-A natural prior, then, would be $\phi(x)=k_{x}/J$.
-Without loss, suppose $\xi_{j}=x$.
-Then, $P(\xi_{j}|\theta,\boldsymbol{\xi})=P(x|\theta)/k_{x}$.
-Thus,
-$\frac{P(j|\textbf{u})}{P(j)}=\frac{P(j|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(\xi_{j}|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(x|\theta)/k_{x}}{1/J}=\frac{P(x|\theta)}{\phi(x)}$
-The need for consumer choice uncertainty.
-Given the non-identifiability problem, it may be tempting to simply assume that a consumer could observe the characteristics of their choices costlessly, and simply use the state-action model to model the uncertainty generated by $\theta$.
-Indeed, this is what most empirical discrete choice models implicitly do when $\xi$ is taken to be the choice variable.
-This is reasonable since the econometrician observes $P(\xi,\theta)$, not $P(k,\textbf{u})$.
-However, there are many circumstances in which the state-characteristic model we propose herein may be helpful.
-For one, the choice probability predicted by the state-action model may be non-unique.
-Second, we may want to incorporate information that some options are more likely than others.
-Third, we may observe choices which appear strictly dominated, as is empirically common [2], which cannot be explained by modeling characteristics as actions in the state-action model.
-Fourth, there is no incorporation of the inherent ubiquity of certain options.
-To illustrate these problems in structural econometric context, suppose Harold Zurcher is buying used buses, and $X=\{\text{red bus, blue bus, lemon bus}\}$ such that $\text{red}=\text{blue}\succ\text{lemon}$ in all states.
-In the state-action model without choice uncertainty, he would ever choose a lemon, no matter how abundant they were.
-And even if we knew red buses outnumbered blue buses 100-to-1, the state-action model does not prescribe a unique optimum policy – it could be optimal for Zurcher to only buy blue (or red).
-So what are we missing?
-What we have shown is that when characteristics are modeled as state variables, the state-action imposes a cost function which is effectively $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$, and when characteristics are modeled as actions, then naturally the cost function is $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$.
-In the former, the DM behaves as though observing $\theta$ is free.
-In the latter, the DM observes $\xi$ for free.
-To wit, the natural extension of the state-action model which we propose in the next section contains two terms: $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$ and $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$.
+In short, the problem with modeling inattention to states and characteristics using the state-action model boils down to the fact that the Shannon information between actions $a$ and random elements $(\omega_{1},\omega_{2})$ does not account for the information between $(a,\omega_{1})$ and $\omega_{2}$.
+In the context of consumer choice over products with arbitrary indices $j\in\{1,...,J\}$, characteristics $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$, and states $\theta$, this is to say the Shannon information between $j$ and $(\boldsymbol{\xi},\theta)$ does not account for information between $\theta$ and $(j,\boldsymbol{\xi})$.
+As a simple example, consider a canonical state-action consumer choice problem.
+We have a health insurance market with two insurance contracts $j\in\{1,2\}$.
+A consumer faces two sources of uncertainty.
+First, he will be high risk ($\theta=HR$) with probability (w.p.) $\frac{1}{2}$ and low risk ($\theta=LR$) w.p.
+$\frac{1}{2}$.
+Second, he knows that one contract has high coverage and the other has low coverage but is uncertain as to how they are permuted: $\boldsymbol{\xi}\equiv(\xi_{1},\xi_{2})=(HC,LC)$ w.p.
+$\frac{1}{2}$ and $\boldsymbol{\xi}\equiv(\xi_{1},\xi_{2})=(LC,HC)$ w.p.
+$\frac{1}{2}$.
+The consumer receives utility 1 if he chooses “correctly” ($HC$ when $\theta=HR$ and $LC$ when $\theta=LR$) and 0 otherwise.
+In the state-action model, the consumer’s information cost from committing to an information policy $Q(j,(\boldsymbol{\xi},\theta))$ is given by the mutual information $I_{Q}$ between the action $j$ and the random element $(\boldsymbol{\xi},\theta)$.
+The cost of an information policy $Q$ which sends the “correct” action recommendation signal with probability $q>\frac{1}{2}$ and the “incorrect” signal with probability 1q1-q is
+$\kappa(Q)=I_{Q}(j,(\boldsymbol{\xi},\theta))=\log 2+q\log q+(1-q)\log(1-q)$ (2)
+where $I_{Q}(\cdot,\cdot)$ is mutual information under probability measure $Q$.
+The problem is that a consumer who already knows his $\theta$ and is only uncertain about $\boldsymbol{\xi}$ also has an optimal information policy that involves receiving the “correct” signal with probability $q$.
+The cost of such a policy is also given by eq.˜2.
+Because the consumer who knows $\theta$ and the consumer who does not both face the exact same objective function, they end up being observationally identical.
+For the consumer who knows $\theta$ already, the state-action information cost function correctly penalizes the information that allows the consumer to choose the “correct” contract with better than random odds – that is, the information gain of $P(\xi|\theta)$ over $\phi(\xi)$, where $\xi$ represents the chosen contract and $\phi(\xi)=\frac{1}{2}$ is the probability of choosing $\xi$ randomly.
+What the information cost $I_{Q}(j,(\boldsymbol{\xi},\theta))$ fails to account for is the information gain in $\theta$.
+This is because when the consumer who does not know $\theta$ receives the signal “choose option j=1j=1” from the oracle, he receives no per se information about his future health risk222in this case, he also receives no information about the contracts – and hence does not have to pay for such information under the state-action cost function.
+The key insight is that although the action recommendation signal $j$ does not provide any information $\theta$ (and $\boldsymbol{\xi}$ does not either), $(j,\boldsymbol{\xi})$ does.
+In our health insurance example, “choose option $j=1$ where $\xi_{1}$ is likely to be HCHCdoes indeed provide information that the state is likely $HR$.
+That is, what we are missing in the state-action cost function is $I_{Q}(\theta,(j,\boldsymbol{\xi}))$ – the mutual information between $\theta$ and $(j,\boldsymbol{\xi})$.
+As it turns out, under the optimal information policy, it suffices to account for the mutual information between $\theta$ and the characteristics $\xi$ of the chosen contract.
+To wit, we present a model which accounts for this.
 Model
-Notation.
-Let $m,m^{\prime}$ be measures defined on the same measurable space.
-$m\ll m^{\prime}$ denotes absolute continuity.
-$m\ll m^{\prime}\ll m$ is interchangeable with $\sim m^{\prime}.
-m=m$$m^{\prime}$ if $m(A)=m^{\prime}(A)$ for each measurable set.
-The KL-divergence between two probabilities is defined
-$D_{KL}(m\|m^{\prime})=\begin{cases}\int\log\left(\frac{dm}{dm^{\prime}}\right)\>dm&(m\ll m^{\prime})\\ \infty&\text{else}\end{cases}$
-By convention, $\infty\cdot 0=0$ for integration purposes.222see [3]
+We now present a model of rational inattention in which information about states and characteristics are both costly.
+While our model is microfounded by a decision problem with discrete actions $j\in A=\{1,...,J\}$ and random elements $(\xi_{1},...,\xi_{J},\theta)\in X^{J}\times\Theta$, the utility of the decision-maker depends only on the state $\theta$ and the chosen characteristic $\xi\equiv\xi_{j}$.
+Thus, we can choose to either model the decision-maker’s information policy over $A\times X^{J}\times\Theta$ or, more directly, over $X\times\Theta$.
+We go with the latter for several reasons.
+The foremost reason is that it is more tractable theoretically.
+But it is also more tractable in applied settings, as econometricians typically have data on the state $\theta$ and the chosen characteristic $\xi$, but not necessarily on $(j,\boldsymbol{\xi},\theta)$.
+Finally, this formulation has the added bonus of being interpretable as an interpolation between two state-action models: one in which $\xi$ is costless to learn, and one in which $\theta$ is costless to learn.
+Despite the tractability, the model is nevertheless abstract: we let the characteristic space and the state space be general Polish spaces.
 Model.
@@
   1. 1.
-    (Bayes plausibility) $P|_{\theta}=\mu$
+    (Bayes plausibility) $P_{\theta}=\mu$
   2. 2.
-    (marginal feasibility) $P|_{\xi}\ll\phi$
-The cost function $\kappa_{\alpha,\lambda}$ can be written in several ways:
+    (marginal feasibility) $P_{\xi}\ll\phi$
+The cost function $\kappa_{\alpha,\lambda}$ is defined
+[…]
+The cost function can also be written equivalently in several equivalent ways:
 […]
@@
     $u:X\times\Theta\to[\underline{u},\overline{u}]$ is real-valued, bounded, and measurable
-Example: state-action with costlessly observed $\xi$.
-When $\alpha=0$, $\kappa_{\alpha,\lambda}(P)$ simplifies to $I_{P}(\xi,\theta)$, which corresponds to the state-action rational inattention model.
-Id est, when $\alpha=0$, the DM can choose characteristics directly, just as in the state-action model, the DM chooses actions directly.
+Microfoundation.
+Consider a decision problem defined by actions $j\in A=\{1,...,J\}$, random elements $(\boldsymbol{\xi},\theta)\in X^{J}\times\Theta$, an exogenous prior $\rho\in\Delta(X^{J}\times\Theta)$, and a utility function $u(\xi,\theta)$ that depends only on the state $\theta$ and the chosen characteristic $\xi$.
+Assume also that:
+  1. 1.
+    (independence) $\boldsymbol{\xi}\bot\!\!\!\bot\theta$
+  2. 2.
+    (a priori homogeneity) the indexing of choices $\{1,...,J\}$ are arbitrary
+  3. 3.
+    (design-based uncertainty) the DM knows the number of choices $\phi(x)$ that correspond to characteristic $x\in X$, but does not know how these characteristics are permuted among the options 1,,J1,...,J
+The DM’s objective is to choose an information policy $P\in\Delta(A\times X^{J}\times\Theta)$ which maximizes expected utility less an information cost subject to Bayes plausibility $P_{X^{J}\times\Theta}=\rho$.
+Where this decision problem departs from the state-action model is that the DM in this decision problem faces an information cost with two terms.
+First, there is the (usual) Shannon information between actions $j$ and random elements $(\boldsymbol{\xi},\theta)$.
+We can re-write this as
+$I_{P}(j,(\boldsymbol{\xi},\theta))$ $=E_{P}\left[\log\left(\frac{P(j,(\boldsymbol{\xi},\theta))}{P(j)\times P((\boldsymbol{\xi},\theta))}\right)\right]$
+$=E_{P}\left[\log\left(\frac{P(j,(\boldsymbol{\xi},\theta)|\theta)}{P(j)\times P(\boldsymbol{\xi}|\theta)}\right)\right]$
+$=E_{P}\left[\log\left(\frac{P(j,(\boldsymbol{\xi},\theta)|\theta)}{P(j|\theta)\times P(\boldsymbol{\xi}|\theta)}\right)\right]$ (homogeneity)
+$=E_{P}\left[\log\left(\frac{P(\boldsymbol{\xi}|j,\theta)}{P(\boldsymbol{\xi}|\theta)}\right)\right]$
+$=E_{P}\left[\log\left(\frac{P(\boldsymbol{\xi}|j,\theta)}{P(\boldsymbol{\xi})}\right)\right]$ (independence)
+Let $\xi_{-j}=(\xi_{1},...,\xi_{j-1},\xi_{j+1},...,\xi_{J})$.
+By complementarity under the permutation assumption, $\xi_{j}$ pins down what characteristics are left in $\xi_{-j}$, but does not pin down how they are permuted.
+If the information policy is optimal, $j$ should provide no additional information about how they are permuted as it is not payoff-relevant by assumption.
+Thus, if $P$ is optimal, then for any $\textbf{x}=(x_{1},...,x_{J})\in X$,
+$\frac{P(\boldsymbol{\xi}=\textbf{x}|j,\theta)}{P(\boldsymbol{\xi}=\textbf{x})}\equiv\frac{P(\boldsymbol{\xi}=(x_{j},x_{-j})|j,\theta)}{P(\boldsymbol{\xi}=(x_{j},x_{-j}))}=\frac{P(\xi=x_{j}|j,\theta)P(\xi_{-j}=x_{-j}|\xi_{j},j,\theta)}{P(\xi_{j}=x_{j})P(\xi_{-j}=x_{-j}|\xi_{j})}=\frac{P(\xi_{j}=x_{j}|j,\theta)}{P(\xi_{j}=x_{j})}\equiv\frac{P(\xi=x_{j}|\theta)}{\phi(x_{j})}$
+where $\xi$ is the chosen characteristic. Hence the first term of the DM’s information cost can be re-written as
+$I_{P}(j,(\boldsymbol{\xi},\theta))=E_{P}\left[\log\left(\frac{P(\xi|\theta)}{\phi(\xi)}\right)\right]=E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$
+The second information cost faced by the DM is for the expected information gain from $P(\theta|j,\boldsymbol{\xi})$ over the exogenous prior $P(\theta)$.
+Again, if the information policy is optimal, then $P(\theta|j,\boldsymbol{\xi})=P(\theta|\xi_{j})$; that is, the chosen $\xi_{j}$ should provide information about $\theta$, but the exact permutation of $\xi_{-j}$ should not, since it is not payoff relevant.
+Hence, this second cost term can be re-written as
+$I_{P}(\theta,(j,\boldsymbol{\xi}))E_{P}[D_{KL}(P(\theta|j,\boldsymbol{\xi})\|P(\theta))]=E_{P}[D_{KL}(P(\theta|\xi)\|P(\theta))]$
+Setting the DM’s cost function to be a weighted sum of these two cost terms yields
+$\kappa_{\alpha,\lambda}(P)$ $=\alpha\lambda I_{P}(j,(\boldsymbol{\xi},\theta))+(1-\alpha)\lambda I_{P}(\theta,(j,\boldsymbol{\xi}))$
+$=\alpha\lambda E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]+(1-\alpha)\lambda E_{P}[D_{KL}(P(\theta|\xi)\|P(\theta))]$
+which is precisely the cost function presented in our model.
+Example: state-action with costlessly observed $\boldsymbol{\xi}$.
+When $\alpha=0$, $\kappa_{\alpha,\lambda}(P)$ simplifies to $I_{P}(\xi,\theta)$, which is exactly the cost function we would get if we treated characteristics as costlessly observed and then applied the state-action model.
 Example: state-action with costlessly observed $\theta$.
-Consider a state-action model in which the DM knows their type $\theta$ a priori.
+Again, consider a state-action microfoundation in which the DM knows their state $\theta$ a priori.
 The DM has an action set $A=\{1,...,J\}$ consisting of $J$ products with $J$ characteristics.
@@
 Denote $\xi$ with the characteristic of the product he eventually chooses.
-The mutual information $I_{P}(j,\boldsymbol{\xi})$ between action $j\in[J]$ and $\boldsymbol{\xi}$ is
-$\int D_{KL}(P(\boldsymbol{\xi}|j)\|P(\boldsymbol{\xi}))\>dP(j)=D_{KL}(P(\boldsymbol{\xi}|j=1)\|P(\boldsymbol{\xi}))=D_{KL}(P(\xi_{1}|j=1)\|P(\xi_{1}))=D_{KL}(P(\xi)\|\phi(\xi))$
+Given $\theta$, the mutual information $I_{P}(j,\boldsymbol{\xi})$ between action $j\in A$ and $\boldsymbol{\xi}$ is
+$\sum_{j\in A}D_{KL}(P(\boldsymbol{\xi}|j,\theta)\|P(\boldsymbol{\xi}))\cdot P(j)=D_{KL}(P(\boldsymbol{\xi}|j=1,\theta)\|P(\boldsymbol{\xi}))=D_{KL}(P(\xi_{1}|j=1,\theta)\|P(\xi_{1}))$
+which is equal to $D_{KL}(P(\xi|\theta)\|\phi(\xi))$.
 The third equality follows from the observation that choosing $j=1$ provides only information about what $\xi_{1}$ and is not (meaning it provides no information about $(\xi_{2},...,\xi_{J})$ beyond the fact that they are not $\xi_{1}$).
 Thus, this corresponds to the state-characteristic model with $\alpha=1$ and $\phi(\xi)=1/J$.
-Example: Maxwell-Boltzmann statistics. When $\alpha=1$, the law of iterated expectations yields
+Connection to Maxwell-Boltzmann statistics. When $\alpha=1$, the law of iterated expectations yields
 $U(P)$ $=E_{\theta\sim\mu}\bigg[E_{P}\bigg[u(\xi,\theta)-\lambda D_{KL}(P(\xi|\theta)\|\phi(\xi))\ |\ \theta\bigg]\bigg]$
@@
 To simplify things, we restrict the DM to choosing disintegrations and densities which satisfy
-$\frac{dP(\xi,\theta)}{d(P|_{\xi}\otimes\mu)(\xi,\theta)}=\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}<\infty$ $\frac{dP(\xi)}{d\phi(\xi)}<\infty$ $\frac{dP(\theta)}{d\mu(\theta)}=1$ (2)
+$\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}=\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}<\infty$ $\frac{dP(\xi)}{d\phi(\xi)}<\infty$ $\frac{dP(\theta)}{d\mu(\theta)}=1$ (4)
 everywhere, not just almost everywhere. The existence of these are established in the following lemmas.
 Lemma 1.
-If $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$, then $P(\cdot|\theta)\ll\phi$ $\mu$-a.s. and $P\ll\phi\otimes\mu$.
+If $P_{\xi}\ll\phi$ and $P_{\theta}=\mu$, then $P(\cdot|\theta)\ll\phi$ $\mu$-a.s. and $P\ll\phi\otimes\mu$.
 Lemma 2.
-Let $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$. There exist valid densities such that eq.˜2 is satisfied.
+Let $P_{\xi}\ll\phi$ and $P_{\theta}=\mu$. There exist valid densities such that eq.˜4 is satisfied.
 The proofs are deferred to the appendix.
 Main Results
-We consider the case where $\alpha\in(0,1)$.
+Because results have been established for $\alpha\in\{0,1\}$, we focus on the case where $\alpha\in(0,1)$.
 We normalize $\lambda$ to $1$, since maximizing $U(P)$ is the same as maximizing $U(P)/\lambda$.
-To obtain results for $\lambda\neq 1$, simply take all results herein in which $u(\xi,\theta)$ and associated additive terms appear and divide them by $\lambda$.
+To obtain results for $\lambda\neq 1$, one can simply take all results herein in which $u(\xi,\theta)$ and associated additive terms appear and divide them by $\lambda$.
+For clarity of exposition, it helps to write out $U(P)$ as a single integral
+$U(P)=\int Y(\xi,\theta;P)\>dP$
 Theorem 3 (Gibbs property).
-If $P$ solves the objective eq.˜1 then
+If $P$ solves the objective eq.˜3 then
 $\xi\mapsto Y(\xi,\theta;P)$ is constant $P(\cdot|\theta)$-a.s. for $\mu$-a.e. $\theta$
 Corollary 3.1 (Weighted multinomial logit).
-If $P$ solves the objective eq.˜1 then
+If $P$ solves the objective eq.˜3 then in the discrete case,
 $P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(\xi,\theta)}}{Z(\theta;P)}$ (5)
@@
 Theorem 4.
-The optimal marginal $P|_{\xi}$ is unique, and therefore the optimal CCP is unique.
+The optimal marginal $P_{\xi}$ is unique, and therefore the optimal CCP is unique.
 Theorem 5.
-If $P$ maximizes $V$, then $P|_{\xi}\ll\phi\ll P|_{\xi}$.
-These follow from corollaries˜7.1, 8 and 15.1. In the discrete case, these can be derived by taking Lagrangians.
+If $P$ maximizes $U$, then $P_{\xi}\ll\phi\ll P_{\xi}$.
+These follow from corollaries˜7.1, 8 and 16.1. In the discrete case, these can be derived by taking Lagrangians.
 Testable restrictions
@@
 For any $\theta$, we have that
-$\frac{P(\xi|\theta)}{P^{\prime}(\xi|\theta)}=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}}{\phi^{\prime}(\xi)^{\alpha}P^{\prime}(\xi)^{1-\alpha}}\frac{Z^{\prime}(\theta;P^{\prime}|_{\xi})}{Z(\theta;P|_{\xi})}$
+$\frac{P(\xi|\theta)}{P^{\prime}(\xi|\theta)}=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}}{\phi^{\prime}(\xi)^{\alpha}P^{\prime}(\xi)^{1-\alpha}}\frac{Z^{\prime}(\theta;P^{\prime}_{X})}{Z(\theta;P_{\xi})}$
 For any pair $\xi_{1},\xi_{2}\in X$, we have
@@
 However, knowing $\phi$ naturally pins down $\alpha$.
-First-Step Orthogonality and Its Implications
+First-Step Orthogonality and the Gibbs Property
+We begin our mathematical results sections by highlighting a few necessary conditions for optimality.
 First-step orthogonality is a property of moments of the form $\int f_{P}(\omega)\>dP(\omega)$, where the integrand $f$ depends on the probability with respect to which we are integrating.
-Loosely, it states that a perturbation in $P$ locally affects the moment only through the re-weighting $dP$.
+Loosely speaking, a moment is said to be first-step orthogonal [^3] This term is borrowed from the double machine learning literature.
+perturbation in $P$ locally affects the moment only through the re-weighting $dP$.
 It constitutes a sufficient condition for the Gibbs property and additive separability.
@@
 ∎
-By the same token if $P$ maximizes $U$ subject to a marginal constraint $P|_{\xi}=\nu$, then $Y$ should be additively separable. More on this later.
+By the same token if $P$ maximizes $U$ subject to a marginal constraint $P_{\xi}=\nu$, then $Y$ should be additively separable. More on this later.
 Lemma 8.
 Let $\alpha>0$. If $P$ maximizes $U$ then the density $\frac{dP(\xi)}{d\phi(\xi)}$ does not get arbitrarily close to 0 and does not become arbitrarily large.
-Schrödinger Potentials
-The problem faced by the DM can be separated into two problems: i) maximizing the objective eq.˜1 subject to a further marginal constraint $P|_{\xi}=\nu$ and ii) solving for the optimal marginal $\nu$.
+Lemma˜8 is a crucial lemma because it guarantees that the optimal marginal $P_{\xi}$ lies on the interior, so it is equally crucial to transparently pinpoint its origin: effectively, it arises from an “Inada condition” on the functional form of the cost of information gain from $P_{\xi}$ to $\phi$.
+To highlight this, consider the discrete case, where the information cost $\kappa_{\alpha}$ can be written as
+$\kappa_{\alpha}$ $=\alpha D_{KL}(P_{\xi}\|\phi)+I_{P}(\xi,\theta)$
+$=\left[\alpha\sum_{\xi\in X}\left(\frac{P(\xi)}{\phi(\xi)}\right)\log\left(\frac{P(\xi)}{\phi(\xi)}\right)\ \phi(\xi)\right]+I_{P}(\xi,\theta)$
+$\equiv\left[\alpha\sum_{\xi\in X}f(\xi)\log(f(\xi))\ \phi(\xi)\right]+I_{P}(\xi,\theta)$
+where $f=P_{\xi}/\phi$ is the likelihood ratio/density.
+For $\alpha>0$ and $\xi$ such that $\phi(\xi)>0$, as $f(\xi)\to 0$, $\frac{\partial\kappa}{df(\xi)}\to$-$\infty$; that is, the marginal cost of reducing $f(\xi)\equiv P(\xi)/\phi(\xi)$ increases to infinity.
+In words, the result that the DM never completely zeroes out the possibility of choosing a lemon ($P(\text{lemon})$) when such lemons exist ($\phi(\text{lemon})>0$) in the market comes from the assumption that as the chance of choosing a lemon goes to 0, it becomes arbitrarily costly to further reduce those chances.
+This is mathematically embodied in this model by the Inada property of $x\log x$.
+Rational Inattention as Entropic Optimal Transport
+The problem faced by the DM can be separated into two problems: i) maximizing the objective eq.˜3 subject to an additional marginal constraint $P_{\xi}=\nu$ and ii) solving for the optimal marginal $\nu$.
 We observe here that the former is an instance of the entropic optimal transport problem.
 The constrained optimal marginal is called the Schrödinger bridge from $\nu$ to $\mu$.
-In lemma˜8, we established that if $P$ maximizes $U$ subject to feasibility and Bayes plausibility, then $P|_{\xi}\sim\phi$ and $dP(\xi)/d\phi(\xi)$ is upper bounded and lower bounded away from zero.
-So in this section, we assume $P|_{\xi}$ is restricted to a marginal $\nu$ for which $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from zero.
-Define ${\cal Q}(\nu)=\{P\in\Delta(X\times\Theta):P|_{\xi}=\nu,P|_{\theta}=\mu\}$. We let
-$V(\nu)=\sup_{P\in{\cal Q}(\nu)}U(P)$
-denote the value of imposing the constraint $P|_{\xi}=\nu$.
+Define $\Pi(\nu,\mu)=\{P\in\Delta(X\times\Theta):P_{\xi}=\nu,P_{\theta}=\mu\}$ to be the set of all joint distributions with marginal restrictions $\nu$, $\mu$.
+Then, we define the value of imposing the second constraint $\nu$ to be
+$V(\nu)=\sup_{P\in\Pi(\nu,\mu)}U(P)$
+denote the value of imposing the constraint $P_{\xi}=\nu$.
 We can rewrite U(P)-U(P) as:
@@
 and turn to minimizing
-$-V(\nu)=\inf_{P\in{\cal Q}(\nu)}\int c(\xi,\theta)\>dP-D_{KL}(P\|\nu\otimes\mu)$ (4)
-where $c(\xi,\theta)=-{u(\xi,\theta)}+\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)\in L^{\infty}(\nu\otimes\mu)$.
-This is called the Entropic Optimal Transport (EOT) problem.
-[^3] For simplicity, we may assume $c(\xi,\theta)$ is totally bounded, since the worse case scenario choice of $\nu$ can be set to $\phi$, in which case c=uc=-u, which we assumed to be totally bounded.
-The following theorems concern properties of EOT and can be found in [6], up to notational changes.
+$-V(\nu)=\inf_{P\in\Pi(\nu,\mu)}\int\underbrace{-u(\xi,\theta)+\alpha D_{KL}(\nu\|\phi)}_{c(\xi,\theta)}\>dP+D_{KL}(P\|\nu\otimes\mu)$ (6)
+where $c$ is uniformly bounded from below by assumption.
+This is the Entropic Optimal Transport (EOT) problem.
+The following theorems apply the key properties of EOT to the state-characteristic model and can be found in [nutz2021introduction], up to notational changes.
+Theorem 9.
+For each $\nu$, there exists a unique Schrödinger bridge $P\in\Pi(\nu,\mu)$ that attains the infimum in eq.˜6.
 Theorem 10 (Existence and uniqueness of potentials).
-If $P$ solves eq.˜4, then there exist functions $a_{\nu}\in L^{1}(\nu),-b_{\nu}\in L^{1}(\mu)$ such that
+Let $P$ be the Schrödinger bridge that solves eq.˜6. There exist functions $a_{\nu}\in L^{1}(\nu),-b_{\nu}\in L^{1}(\mu)$ such that
 $\frac{dP}{d(\nu\otimes\mu)}=e^{-a_{\nu}(\xi)-b_{\nu}(\theta)-c(\xi,\theta)}=e^{u(\xi,\theta)-a_{\nu}(\xi)-b_{\nu}(\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}$ (7)
 These are called the Schrödinger potentials.
-Conversely, if there exists potentials such that eq.˜5 holds, then $P$ must be optimal.
+Conversely, if there exist potentials such that eq.˜7 holds, then $P$ must be optimal.
 Moreover, potentials are unique up to translation.
 [^4] That is, for any two pairs of potentials $(a_{\nu},b_{\nu})$ and $(a_{\nu}^{\prime},b_{\nu}^{\prime})$, we have $a_{\nu}-a^{\prime}_{\nu}=b_{\nu}^{\prime}-b_{\nu}$.
+The statement above says that $a_{\nu}$ is in $L^{1}(\nu)$; but because $\nu$ is endogenous, we would much rather have a statement that says $a_{\nu}\in L^{1}(\phi)$.
+We established in lemma˜8 that if $\alpha>0$, then $\frac{d\nu}{d\phi}$ is bounded away from 0 and $\infty$.
+Thus, a stronger statement holds:
 Corollary 10.1.
-If $P$ solves eq.˜4 then $Y(\xi,\theta;P)=a_{\nu}(\xi)+b_{\nu}(\theta)$ and so $U(P)=E_{\nu}[a_{\nu}(\xi)]+E_{\mu}[b_{\nu}(\theta)]$
+Suppose $\frac{d\nu}{d\phi}$ is bounded away from 0 and $\infty$.
+Then, $f\in L^{1}(\nu)\Leftrightarrow f\in L^{1}(\phi)$.
+$P$ solves eq.˜6 if and only if there exist functions $a_{\nu}\in L^{1}(\phi),-b_{\nu}\in L^{1}(\mu)$ satisfying eq.˜7.
+Proof.
+Let $f\in L^{1}(\nu)$. Then
+$\bigg\|\frac{d\nu}{d\phi}\bigg\|_{\infty}\int|f|\ d\nu\geq\int|f|\ \frac{d\nu}{d\phi}\ d\phi=\int|f|\ d\phi$
+Conversely, if $f\in L^{1}(\phi)$ then
+$\bigg\|\frac{d\phi}{d\nu}\bigg\|_{\infty}\int|f|\ d\phi\geq\int|f|\ \frac{d\phi}{d\nu}\ d\nu=\int|f|\ d\nu$
+A trivial consequence is additive separability of the integrand when $P$ is optimal.
+Corollary 10.2.
+If $P$ solves eq.˜6 then $Y(\xi,\theta;P)=a_{\nu}(\xi)+b_{\nu}(\theta)$ and so $U(P)=E_{\nu}[a_{\nu}(\xi)]+E_{\mu}[b_{\nu}(\theta)]$
+Two results from EOT theory prove useful for studying this system.
+The first are the Schrödinger Equations which characterise the relationship between the Schrödinger potentials.
+The second is the duality formula for EOT.
 Theorem 11 (Schrödinger Equations).
@@
 Ergo,
-$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}}\>d\nu(\xi)}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{b_{\nu}(\theta)}}\>d\mu(\theta)}\>d\nu(\xi)$
+$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi^{\prime},\theta)}}{e^{a_{\nu}(\xi^{\prime})}\left(\frac{d\nu(\xi^{\prime})}{d\phi(\xi^{\prime})}\right)^{\alpha}}\>d\nu(\xi)}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta^{\prime})}}{e^{b_{\nu}(\theta^{\prime})}}\>d\mu(\theta^{\prime})}\>d\nu(\xi)$
+If $P$ is optimal and $\nu=P_{\xi}$ is the marginal, then the potential $a_{\nu}(\xi)$ should be constant, since the marginal for $X$ is a choice variable.
+After normalizing $a_{\nu}=0$, it should be obvious that the potential $b_{\nu}(\theta)$ equals the log of the partition function:
+$b_{\nu}(\theta)=\log(Z(\theta;P))$
 Theorem 12 (Duality).
-$V(\nu):=\sup_{P:\ P|_{\xi}=\nu,P|_{\theta}=\mu}U(P)$
+$V(\nu):=\sup_{P\in\Pi(\nu,\mu)}U(P)$
 $V(\nu)$ $=\inf_{a\in L^{1}(\nu),\ b\in L^{1}(\mu)}\int a\>d\nu+\int b\>d\mu+\iint e^{-a-b-c}\>d\nu\>d\mu-1$
 $=\inf_{a\in L^{1}(\nu),\ b\in L^{1}(\mu)}\int a\>d\nu+\int b\>d\mu+\iint e^{u(\xi,\theta)-a-b}\>d\phi^{\alpha}d\nu^{1-\alpha}\>d\mu-1$
-Note that because $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from 0, by Hölder’s equality with coefficients $p=1$ and $q=\infty$, we can interchange $a\in L^{1}(\nu)$ with $a\in L^{1}(\phi)$.
-Corollary 11.1.
-Consider a path $\nu_{\varepsilon}$ such that at $\varepsilon=0$, $\varepsilon\mapsto\frac{d\nu_{\varepsilon}(\xi)}{d\phi(\xi)}$ is continuous for $\phi$-a.e.
-$\xi$, and $d\nu_{\varepsilon}/d\phi$ is uniformly bounded away from 0 and $\infty$ on $\varepsilon\in(-r,r)$.
-Then $\varepsilon\mapsto V(\nu_{\varepsilon})$ is continuous at 0.
-In the discrete case, the Schrödinger potentials can be viewed as the Lagrange multipliers for the marginal constraints.
-In other words, $a_{\nu}(x)$ can be viewed as the rate of increase in $V$ when adding an infinitessimal mass of $x$ to $\nu$.
-This is because
-$V(\nu)=\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
-so in the discrete case, applying the Envelope theorem and normalizing $\int a_{\nu}\>d\nu=0$ gives
-$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}V\bigg{(}(1-\varepsilon)\nu+\varepsilon\delta_{x}\bigg{)}=a_{\nu}(x)$
-The continuous case is more involved and discussion is deferred to the appendix.
 The Jensen Upper Bound of the Marginal Value
-If one writes out the objective – a function of the joint $P$ – swaps the log with the first integral, one obtains a function of the marginal $P|_{\xi}$.
-When $\alpha\in(0,1)$ this function is strictly convex, and, by Jensen’s Inequality, upper bounds $V$.
-We show that their maxima coincide, meaning that solving for the optimal marginal is a relatively simple convex optimization problem.
-Define $f(\nu)=E_{\theta\sim\mu}\left[\log\left(E_{\xi\sim\phi}\left[e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}\right]\right)\right]\equiv E_{\theta\sim\mu}[\log(Z(\theta;\nu)]$
+If one writes out the objective – a function of the joint $P$ – swaps the log with the first integral, one obtains a function of the marginal.
+When $\alpha\in(0,1)$ this function is strictly concave, and, by Jensen’s Inequality, upper bounds $V$.
+We show that their maxima coincide, meaning that solving for the optimal marginal is a relatively simple optimization problem.
+Define
+$f(\nu)=E_{\theta\sim\mu}\left[\log\left(E_{\xi\sim\phi}\left[e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}\right]\right)\right]\equiv E_{\theta\sim\mu}[\log(Z(\theta;\nu))]$
+and, as before,
+$V(\nu)\equiv\sup_{P\in\Pi(\nu,\mu)}U(P)=\sup_{P\in\Pi(\nu,\mu)}E_{(\xi,\theta)\sim P}\left[\log\left(e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\left(\frac{dP(\xi,\theta)}{d(\nu\otimes\mu)(\xi,\theta)}\right)^{-1}\right)\right]$
+The primary claim of this section is that
+$\sup_{\nu\ll\phi}f(\nu)=\sup_{\nu\ll\phi}V(\nu)$
+The secondary claim is that $f$ is strictly concave. These two claims imply that it suffices to maximize the strictly concave function $f(\nu)$.
+Theorem˜13 is an application of Jensen’s inequality, establishing that $f\geq V$.
+Theorem˜14 establishes that $f$ strictly concave and hence is uniquely maximized.
+It follows that to show $V$ is uniquely maximized, it suffices to show that the maximum of $f$ and $V$ coincide, as shown in theorem˜16.
+The immediate corollary˜16.1 is that $V$ is uniquely maximized, and therefore $U$ is uniquely maximized.
 Theorem 13.
-$f(P|_{\xi})\geq U(P)$
+$f(P_{\xi})\geq U(P)$
 Theorem 14.
@@
 Theorem 15.
-If $P$ maximizes $U$ then $f(P|_{\xi})=U(P)$.
+If $P$ maximizes $U$ subject to Bayes plausibility, then $f(P_{\xi})=U(P)$. Hence, $f(P_{\xi})=\sup_{\nu\ll\phi}V(\nu)$
 Proof.
@@
 Theorem 16.
-The maximum of $f$ and $U$ coincide. That is, if $\nu$ maximizes $f$, then there exists a $P$ with $P|_{\xi}=\nu$ which maximizes $U$.
+The maximum of $f$ and $V$ coincide. That is, if $\nu$ maximizes $f$, then there exists a $P$ with $P_{\xi}=\nu$ which maximizes $U$.
+Because each marginal $\nu$ admits a unique Schrödinger bridge $P\in\Pi(\nu,\mu)$, we arrive at our main result:
 Corollary 16.1.
 The maximum of $U$ is unique.
+Discussion
+Schrödinger potentials can be interpreted as the infinite-dimensional analog to the Lagrange multiplier.
+The Schrödinger potential $a_{\nu}(\xi)$ represents a shadow “probability price” on $\nu(\xi)$; it follows that if $\nu$ is optimal, then $a_{\nu}(\xi)$ should be constant a.e.
+Plugging this fact into the Schrödinger equations, it follows that the optimal marginal $\nu$ satisfies
+$\int\frac{e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}}{\int e^{u(\xi^{\prime},\theta)}\left(\frac{d\nu(\xi^{\prime})}{d\phi(\xi^{\prime})}\right)^{1-\alpha}\>d\phi(\xi^{\prime})}\>-1\ d\mu(\theta)=0$ (8)
+This is a state-characteristic analog of the necessary and sufficient condition derived by [caplin2019rational].
+[^5] Recall that theirs involved an inequality because when $\alpha=0$, there can be “corner solutions.” From staring at the definition of $f$, one can glean that eq.˜8 is a first-order condition for $f$ to be maximized on the interior.
+[^6] If one needs any convincing, set $X$ to be discrete, and write out the first-order condition corresponding to a Gateaux derivative perturbation in the direction of a Dirac probability mass $\delta_{x}$.
+As theorem˜14 showed, $f$ is strictly concave on a convex set, implying a unique point satisfying this first-order condition, if it exists.
+It is then left to show that the maximum of $f$ likewise does not lie on the boundary.
+Computationally, eq.˜8 implies that finding the optimal $\nu$ amounts to finding the fixed-point of
+$Tg(\xi)=\int\frac{e^{u(\xi,\theta)}g(\xi)^{1-\alpha}}{\int e^{u(\xi^{\prime},\theta)}g(\xi^{\prime})^{1-\alpha}d\phi(\xi^{\prime})}\>d\mu(\theta)$
+In the model, our goal was to maximize $U$ subject to one marginal constraint $P_{\theta}=\mu$.
+But since we have taken this unusual approach of first maximizing subject to two constraints $P\in\Pi(\nu,\mu)$ and then optimizing the first constraint $\nu$, it makes ample sense to consider the marginal benefit of relaxing our constraints.
+In the discrete case, the Schrödinger potentials can be directly interpreted as the Lagrange multipliers for the marginal constraints.
+In other words, $a_{\nu}(x)$ can be viewed as the rate of increase in $V$ when transfering an infinitessimal mass away from all other elements of $X$ to $\nu(x)$.
+This is because
+$V(\nu)=\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
+via duality, so in the discrete case, applying the Envelope theorem and normalizing $\int a_{\nu}\>d\nu=0$ gives
+$\frac{d}{d\varepsilon}\bigg|_{\varepsilon=0}V\bigg((1-\varepsilon)\nu+\varepsilon\delta_{x}\bigg)=a_{\nu}(x)-\displaystyle\int a_{\nu}\>d\nu=a_{\nu}(x)$
+where $\delta_{x}$ is the Dirac probability at $x$. The continuous case is more involved and discussion is deferred to the appendix. For now, let us assume the premise is true.
+To consider the marginal benefit of relaxing a constraint, consider a concrete example of a DM choosing an insurance policy.
+Utility depends on consumption, which at time 0 is given as $c_{0}$ and at time 1 is given as $c_{1}(\xi,\theta)$ – depending on insurance contract $\xi$ and state $\theta$.
+Suppose we restricted the DM to choosing information policies $P\in\Pi(\nu,\theta)$, where $\nu$ is sub-optimal.
+Then the marginal benefit in utils of shifting $\nu$ infinitessimally towards $\delta_{x}$ is $a_{\nu}(x)$.
+Using logarithmic utility (which makes sense since we assume entropic costs) for time 0 consumption, i.e.,
+$V_{0}(\nu)=\log c_{0}-\alpha D_{KL}(\nu\|\mu)+\sup_{P\in\Pi(\nu,\mu)}\iint u(c_{1}(\xi,\theta))\>dP(\xi,\theta)-I_{P}(\xi,\theta)$
+we could then use marginal utility of consumption $\partial V_{0}/\partial c_{0}=1/c_{0}$ as a welfare numeraire.
+$a_{\nu}(x)\cdot c_{0}$ can be interpreted as the “probability price” in terms of time 0 consumption – the willingness to pay – to shift $\nu$ marginally towards $\delta_{x}$.
+Likewise, if we instead normalized $b_{\nu}$, then $b_{\nu}(\theta)$ can be viewed as the “probability price” of changing the underlying probability distribution over states of the world to add additional probability to state $\theta$.
+Because $\mu$ is generally not optimal – the DM takes it as given – this is more of a hypothetical.
+However, it could be useful for considering the welfare effect of actual shifts in probability – for example, the welfare change from decreasing the probability of a certain risk, given that consumers are rationally inattentively insured.
+Computation
+We turn to the computational problem of finding the optimal joint distribution $P$ which solves eq.˜3 when $X$ and $\Theta$ are finite sets.
+By eq.˜5, it suffices to solve for the optimal marginal $P_{\xi}$ to obtain the optimal conditional choice probability.
+We suggest three algorithms for doing so.
+Fixed-Point Iteration. [caplin2019rational] suggest the Blahut-Arimoto (BA) algorithm as a method for computing the optimal value. The direct analogue of BA is
+$g_{t+1}(\xi)=[Tg_{t}](\xi)=\int\frac{e^{u(\xi,\theta)}g_{t}(\xi)^{1-\alpha}}{\int e^{u(\xi^{\prime},\theta)}g_{t}(\xi^{\prime})^{1-\alpha}d\phi(\xi^{\prime})}\>d\mu(\theta)$ $g_{t}(\xi)=\frac{d\nu_{t}(\xi)}{d\phi(\xi)}$
+and takes advantage of the fact that $g$ corresponds to the optimal $\nu$ if and only if $Tg=g$. Fixed-point iteration converges nicely for the majority of reasonably conceivable problems.
+Convex optimization.
+An alternative to fixed-point iteration is convex optimization of f-f.
+The results establish f-f is a strictly convex function over the convex set $\Delta X$ which is known to be minimized on $\text{int}(\Delta X)$, and that the $\nu$ which minimizes f-f is the $\nu$ which maximizes $V$.
+Numerical convex optimization methods like mirror ascent work reasonably well but can scale poorly.
+The main pitfall of this is that if the optimal $\nu$ lies near the boundary, the slope between $\nu$ and the boundary is extremely steep.
+Sinkhorn’s Algorithm.
+For discrete $\xi,\theta$, the problem of finding Schrödinger potentials such that eq.˜7 holds is equivalent to the matrix scaling problem.
+Formally, we can re-write eq.˜7 in the discrete case as
+$P(\xi,\theta)=e^{-a_{\nu}(\xi)}\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]e^{-b_{\nu}(\theta)}$ (9)
+If we think of the middle term as a $|X|\times|\Theta|$ matrix $A$ and $P$ as a $|X|\times|\Theta|$ matrix such that the marginals are $\nu$ and $\mu$ respectively, we can re-formulate the problem as follows: given a matrix A and a pair of probability vectors $\boldsymbol{\nu}$, $\boldsymbol{\mu}$, find diagonal matrices $\textbf{D}_{1},\textbf{D}_{2}$ such that for $\textbf{P}:=\textbf{D}_{1}\textbf{A}\textbf{D}_{2}$ is a probability matrix whose columns sum to $\sum_{\xi}P_{\xi\theta}=\mu_{\theta}$ and rows sum to $\sum_{\theta}P_{\xi\theta}=\nu_{\xi}$.
+For our purposes, A is the matrix whose $(\xi,\theta)$-entry is $\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]$.
+The matrix scaling problem has a known solution: Sinkhorn’s algorithm.
+For a given A, Sinkhorn’s algorithm solves for the optimal potentials $a_{\nu}$, $b_{\nu}$, which allow us to back out the optimal $P\in\Pi(\nu,\mu)$ via eq.˜9.
+Theorem 17 (Sinkhorn’s algorithm).
+[nutz2021introduction, Theorem 6.15] Consider the EOT problem. Set $a_{t}=0$. For $t\geq 0$, set
+$b_{t}(\theta)=\log\left(\int e^{u(\xi,\theta)-a_{t}(\xi)}\>d\phi^{\alpha}\>d\nu^{1-\alpha}\right)$
+$a_{t+1}(\xi)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)$
+We may define $P_{2t}$ by via the potentials $a_{t}$, $b_{t}$, and $P_{2t-1}$ via the potentials $a_{t},b_{t-1}$.
+Then:
+  1. 1.
+    $a_{t}$ and $b_{t}$ converge pointwise and in their respective $L^{p}$-norms, $1\leq p<\infty$
+  2. 2.
+    $D_{KL}(P\|P_{t})\to 0$
+  3. 3.
+    $D_{KL}(P_{t}\|R)\to D_{KL}(P\|R)$
+  4. 4.
+    $P_{t}\to P$ in total variation
+Although Sinkhorn’s algorithm allows us to find the Schrödinger potentials for a given $\nu$, it does not tell us what the optimal $\nu$ should be.
+However, observe that the Schrödinger potential $a_{t}(x)$, after normalization, tells us how much $V$ improves when we update $\nu$ to add mass to $x$.
+It can be thought of as the “direction of steepest ascent,” and so it makes sense to integrate a “gradient ascent” step into the algorithm which does not even require computing a gradient (since we already have $a_{t}$ from Sinkhorn iteration).
+One such procedure is given by
+$b_{t}(\theta)=\log\left(\int e^{u(\xi,\theta)-a_{t}(\xi)}\>d\phi^{\alpha}\>d\nu_{t}^{1-\alpha}\right)$
+$a_{t+1}(\xi)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)-\alpha\log\left(\frac{d\nu_{t}(\xi)}{d\phi(\xi)}\right)$
+$d\nu_{t+1}(\xi)=\frac{d\nu_{t}(\xi)\exp\left(\eta_{t+1}a_{t+1}(\xi)\right)}{\int\exp\left(\eta_{t+1}a_{t+1}(\xi^{\prime})\right)d\nu_{t}(\xi^{\prime})}$
+where $\eta_{t}$ is tunable. The update on $\nu_{t+1}$ is equivalent to
+$\log\left(\frac{d\nu_{t+1}(\xi)}{d\phi(\xi)}\right)\propto\log\left(\frac{d\nu_{t}(\xi)}{d\phi(\xi)}\right)+\eta_{t+1}\left(a_{t+1}(\xi)-E_{\phi}[a_{t+1}]\right)$
+The above-specified algorithm updates $\nu$ every time $a$ and $b$ are updated, but this can also be tuned (i.e. to allow for several updates of $a$ and $b$ for every update of $\nu$).
+This procedure can be faster than the other methods, especially for large $|X|\times|\Theta|$.
+Generally, it works better when the optimal $\nu$ lies farther from the boundary, which is typically the case when the cost of acquiring information about $\xi$ – i.e.
+$\alpha\lambda$ – is not too low.
 Conclusion
@@
 Let $P(\cdot|\theta)$ be any disintegration kernel. To see that $P(\cdot|\theta)\ll\phi$, $\mu$-a.s., let $\phi(A)=0$. Then,
-$P|_{\xi}(A)=\int P(A|\theta)\>d\mu(\theta)=0$
+$P_{\xi}(A)=\int P(A|\theta)\>d\mu(\theta)=0$
 which implies that $P(A|\theta)=0$ $\theta$-a.s. Therefore, $dP(\xi|\theta)/d\phi(\xi)<\infty$ $\phi\otimes\mu$-a.e. To see $P\ll\phi\otimes\mu$, let $(\phi\otimes\mu)(S)=0$. Then,
@@
 $\int\frac{d}{d\varepsilon}\bigg|_{\varepsilon=t^{+}}\frac{dP_{\varepsilon}(\xi)}{d\phi(\xi)}\log\left(\frac{dP_{\varepsilon}(\xi)}{d\phi(\xi)}\right)\>d\phi(\xi)=0$
-The exchangeability of the integral and the derivative follows from [3, Theorem 2.27]; in particular, the sufficient condition is that the absolute value of the derivative of the integrand is bounded by an element of $L^{1}$.
+The exchangeability of the integral and the derivative follows from [folland1999real, Theorem 2.27]; in particular, the sufficient condition is that the absolute value of the derivative of the integrand is bounded by an element of $L^{1}$.
 The result follows by construction of $Y$.
@@
 so $e^{b(\theta)}=\infty$ everywhere, which also raises a contradiction. ∎
-Proof of corollary˜11.1.
-Suppose $\gamma,\nu$ are two marginals, and without loss $V(\gamma)\geq V(\nu)$.
-We can upper bound $V(\gamma)$ by using the “wrong” Schrödinger potentials in the dual objective.
-In particular, let $\gamma=\nu_{\varepsilon}$
-$|V(\nu_{\varepsilon})-V(\nu)|\leq\bigg{|}\int a_{\nu}\>d(\nu_{\varepsilon}-\nu)+\iint e^{u(\xi,\theta)-a_{\nu}-b_{\nu}}\left[\left(\frac{d\nu_{\varepsilon}}{d\phi}\right)^{1-\alpha}-\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\right]\>d\phi\>d\mu-1\bigg{|}$
-But this is continuous w.r.t.
-$\varepsilon$.
-Since $\frac{d\nu_{\varepsilon}}{d\phi}$ is uniformly bounded away from 0 and $\infty$ in a neighbourhood of $\varepsilon=0$, we apply the Dominated Convergence Theorem to get $V(\nu_{\varepsilon})\to V(\nu)$.
-∎
-Proof of theorem˜12.
+Proof of theorem˜13.
 By Jensen’s inequality,
@@
 $=E_{\theta\sim\mu}\left[\log\left(\int e^{u(\xi,\theta)}\left(\frac{dP(\xi)}{d\phi(\xi)}\right)^{1-\alpha}d\phi(\xi)\right)\right]$ (change of measure)
-$=f(P|_{\xi})$
-Proof of theorem˜13.
+$=f(P_{\xi})$
+Proof of theorem˜14.
 Let $\gamma,\nu$ be feasible probability measures and $\beta\in(0,1)$. Since $\alpha\in(0,1)$, $x\mapsto x^{1-\alpha}$ is strictly concave, it follows
@@
 as needed. ∎
-Proof of theorem˜14.
-If $\nu$ maximizes $f$, then a first-order condition is that for any $\gamma$ which defines a path $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$, we have $df(\nu_{\varepsilon})/d\varepsilon\leq 0$ at 0, which implies:
+Proof of theorem˜16.
+If $\nu$ maximizes $f$, then a first-order condition is that for any $\gamma\ll\phi$ which defines a path $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$, we have $df(\nu_{\varepsilon})/d\varepsilon\leq 0$ at 0, which implies:
 $0$ $\geq\frac{df(\nu_{\varepsilon})}{d\varepsilon}\bigg|_{\varepsilon=0}\geq\int\frac{\int(1-\alpha)e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\frac{d(\gamma-\nu)(\xi)}{d\phi(\xi)}d\phi(\xi)}{\int e^{u(\xi,\theta)}d\phi^{\alpha}d\nu^{1-\alpha}}d\mu=\int\frac{\int e^{u(x,\theta)}\left(\frac{d\nu(x)}{d\phi(x)}\right)^{-\alpha}\>d\gamma(x)}{\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\>d\nu(\xi)}-1\>d\mu$
@@
 $dP(\xi,\theta)=\frac{e^{u(\xi,\theta)}d\phi(\xi)^{\alpha}d\nu(\xi)^{1-\alpha}}{e^{b(\theta)}}$
-is not only a valid probability measure, but also $P|_{\xi}=\nu$, $P|_{\theta}=\mu$, and $P$ maximizes $U$ subject to those constraints. ∎
-Proof of corollary˜15.1.
-If $P,P^{\prime}$ both maximize $U$, then $P|_{\xi},P^{\prime}|_{\xi}$ both maximize $f$.
-Since $f$ is strictly concave, it follows that the marginals are equal: $P|_{\xi}=P^{\prime}|_{\xi}=\nu$.
+is not only a valid probability measure, but also $P_{\xi}=\nu$, $P_{\theta}=\mu$, and $P$ maximizes $U$ subject to those constraints. ∎
+Proof of corollary˜16.1.
+If $P,P^{\prime}$ both maximize $U$, then $P_{\xi},P^{\prime}_{X}$ both maximize $f$.
+Since $f$ is strictly concave, it follows that the marginals are equal: $P_{\xi}=P^{\prime}_{X}=\nu$.
 Because the joint must satisfy the MNL formula, it follows $P$ and $P^{\prime}$ are just versions of one another.
@@
 The Envelope
-As suggested, the Envelope Theorem provides a nice heuristic interpretation for $a_{\nu}(x)$.
-It is left to show, however, that $V(\nu_{\varepsilon})$ or $V(\nu_{\varepsilon})$ can be written as the supremum of a function whose derivative is equicontinuous [5].
+As suggested, the Envelope Theorem provides a nice heuristic interpretation for the potential $a_{\nu}(x)$.
+We have left to show, however, that $V(\nu_{\varepsilon})$ or $V(\nu_{\varepsilon})$ can be written as the supremum of a function whose derivative is equicontinuous [milgrom2002envelope].
 If $b_{\nu}$ is a Schrödinger potential, then plugging in $a_{\nu}=\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b_{\nu}}\>d\mu\right)$ gives
@@
 $V(\nu)=-\alpha D_{KL}(\nu\|\phi)+\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)\>d\nu$
-By construction of $V(\nu)=\sup_{P\in{\cal Q}(\nu)}U(P)$,
+By construction of $V(\nu)=\sup_{P\in\Pi(\nu,\mu)}U(P)$,
 $\overline{u}-\alpha D_{KL}(\nu\|\phi)\geq V(\nu)$
@@
 Define $g:B\times(-r,r)\to\mathbb{R}$ by
-$g(b,\varepsilon)=\alpha D_{KL}(\nu\|\phi)+\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu$
+$g(b,\varepsilon)=\alpha D_{KL}(\nu_{\varepsilon}\|\phi)+\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu_{\varepsilon}$
 Then,
@@
 $\frac{dV(\nu_{\varepsilon,n})}{d\varepsilon}\bigg|_{\varepsilon=0}\to_{n}\int a_{\nu}\>d(\delta_{x}-\nu)=a_{\nu}(x)-\int a_{\nu}\>d\nu$
-Sinkhorn’s algorithm
-In the discrete case, the problem of finding Schrödinger potentials such that eq.˜5 holds is equivalent to the matrix scaling problem. Formally, we can re-write eq.˜5 in the discrete case as
-$P(\xi,\theta)=e^{-a_{\nu}(\xi)}\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]e^{-b_{\nu}(\theta)}$
-If we think of the middle term as a $|X|\times|\Theta|$ matrix $A$, the Schrödinger potentials as diagonal matrices, and $P$ as a $|X|\times|\Theta|$ matrix such that the marginals are $\nu$ and $\mu$ respectively, we can re-formulate the problem as follows: given a matrix A and a pair of probability vectors $\boldsymbol{\nu}$, $\boldsymbol{\mu}$, find diagonal matrices $\textbf{D}_{1},\textbf{D}_{2}$ such that for $\textbf{P}:=\textbf{D}_{1}\textbf{A}\textbf{D}_{2}$ is a probability matrix whose columns sum to $\sum_{\xi}P_{\xi\theta}=\mu_{\theta}$ and rows sum to $\sum_{\theta}P_{\xi\theta}=\nu_{\xi}$.
-Theorem 16 (Sinkhorn’s algorithm).
-Set $a_{t}=0$. For $t\geq 0$, set
-$b_{t}(\theta)=\log\left(\int e^{u(\xi,\theta)-a_{t}(\xi)}\>d\phi^{\alpha}\>d\nu^{1-\alpha}\right)$
-$a_{t+1}(\theta)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)$
-We may define $P_{2t}$ by via the potentials $a_{t}$, $b_{t}$, and $P_{2t-1}$ via the potentials $a_{t},b_{t-1}$.
-Then:
-  1. 1.
-    $a_{t}$ and $b_{t}$ converge pointwise and in their respective $L^{p}$-norms, $1\leq p<\infty$
-  2. 2.
-    $D_{KL}(P\|P_{t})\to 0$
-  3. 3.
-    $D_{KL}(P_{t}\|R)\to D_{KL}(P\|R)$
-  4. 4.
-    $P_{t}\to P$ in total variation
+Concavity of $V$
+Intuition via duality. Let $g(\nu)$ be
+$g(\nu)=-\sup_{P\in\Pi(\mu,\nu)}\iint\log\left(e^{u(\xi,\theta)}\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
+Then, the dual is
+$g^{*}(\psi)$ $=\sup_{\nu}\int\psi\>d\nu+\sup_{P\in\Pi(\mu,\nu)}\iint\log\left(e^{u(\xi,\theta)}\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
+$=\sup_{\nu}\sup_{P\in\Pi(\mu,\nu)}\iint\log\left(e^{u(\xi,\theta)+\psi(\xi)}\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
+$=\sup_{\nu}\int\psi\>d\nu+\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
+where $\overline{\nu}\equiv\overline{\nu}_{\psi},\overline{P}\equiv\overline{P}_{\psi}$ denote the optimal choices of $\nu,P$ given $\psi$.
+Consider a path $\psi_{\varepsilon}=(1-\varepsilon)\psi+\varepsilon H$.
+Even though the optimal choice $\overline{\nu}$ depends on $\psi$, by the Envelope Theorem,
+$\frac{dg^{*}(\psi_{\varepsilon})}{d\varepsilon}=\int H-\psi\>d\overline{\nu}$
+Now consider the biconjugate
+$g^{**}(\nu)=\sup_{\psi}\int\psi\>d\nu-g^{*}(\psi)$
+For $\psi$ to be chosen optimally, it must be the case that for any direction $H$,
+$\int(H-\psi)\>d(\nu-\overline{\nu}_{\psi})=0$
+from which it follows that $\psi$ is chosen such that $\nu=\overline{\nu}_{\psi}$. With this being the case,
+$g^{**}(\nu)=-\int a_{\nu}\>d\nu-\int b_{\nu}\>d\mu=-V(\nu)$
+An Approach Without First-Order Conditions
+Consider first the outer integrand of $f$, given $\nu$:
+$f(\nu)=\int\Lambda_{\nu}(\theta)\>d\mu(\theta)$
+$\Lambda_{\nu}(\theta)=\log\left(\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}d\phi(\xi)\right)=\log\left(\int\exp\left(\underbrace{u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)}_{h_{\nu,\theta}(\xi)}\right)\ d\nu(\xi)\right)$
+The Donsker-Varadhan variational formula states that a log-sup-exp is, in fact, a value function of sorts:
+$\Lambda_{\nu}(\theta)=\log\left(\int\exp(h_{\theta}(\xi))\>d\phi(\xi)\right)$ $=\sup_{Q_{\theta}\ll\phi}\bigg\{E_{Q}[h_{\nu,\theta}]-D_{KL}(Q_{\theta}\|\nu)\bigg\}$
+$=\sup_{Q_{\theta}\ll\phi}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dQ_{\theta}(\xi)}{d\nu(\xi)}\right)\>dQ_{\theta}(\xi)\bigg\}$
+Then,
+$f(\nu)=\int\sup_{Q_{\theta}\ll\phi}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dQ_{\theta}(\xi)}{d\nu(\xi)}\right)\>dQ_{\theta}(\xi)\bigg\}\>d\mu(\theta)$
+Under suitable regularity conditions, Rockafellar’s interchange theorem allows for the exchange of $\sup$ and $\int$. Assuming such conditions hold,
+$f(\nu)=\sup_{Q_{\theta}\ll\phi}\bigg\{\iint u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dQ_{\theta}(\xi)}{d\nu(\xi)}\right)\>dQ_{\theta}(\xi)\>d\mu(\theta)\bigg\}$
+Define $P(\xi,\theta)=Q_{\theta}(\xi)\otimes\mu(\theta)$. Then, we have
+$f(\nu)$ $=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(\nu\otimes\mu)(\xi,\theta)}\right)\>dP(\xi,\theta)\bigg\}$
+Lastly, how do we know that $P_{\xi}=\nu$?
+We don’t: because only at the supremum is this true.
+But to see that this is true at the supremum, decompose the RHS so that $P_{\xi}$ is plugged in everywhere $\nu$ is:
+$f(\nu)$ $=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{dP_{\xi}(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}\right)+(1-\alpha)\log\left(\frac{d\nu(\xi)}{dP_{\xi}(\xi)}\right)\>dP(\xi,\theta)\bigg\}$
+$=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{dP_{\xi}(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}\right)\>dP(\xi,\theta)-(1-\alpha)D_{KL}(P_{\xi}\|\nu)\bigg\}$
+And so
+$\sup_{\nu\ll\phi}f(\nu)$ $=\sup_{P:P_{\xi}\ll\phi,\ P_{\theta}=\mu}\bigg\{\int u(\xi,\theta)-\alpha\log\left(\frac{dP_{\xi}(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}\right)\>dP(\xi,\theta)\bigg\}$
+$=\sup_{\nu\ll\phi}V(\nu)$
+which is precisely the DM’s problem.
+As a Selection Device for State-Action Rational Inattention
+It is a natural to use the state-characteristic model as a selection device when state-action yields non-unique solutions.
+Let $V_{\alpha}$ denote the value function of the marginal given $\alpha$, and let $v^{*}_{\alpha}=\arg\max V_{\alpha}$, for $\alpha\in(0,1)$.
+If $\arg\max V_{0}$ is not a singleton – i.e.
+multiple solutions in the state-action model – then one can use the state-characteristic model as a selection device by setting $\nu_{0}^{*}=\lim_{\alpha\to 0}\nu^{*}_{\alpha}$.
+Since the optimal marginal $\nu^{*}_{\alpha}$ is unique for each $\alpha$.
+If such an limit exists, it turns out to be the minimum divergence marginal:
+$\lim_{\alpha\to 0}\nu^{*}_{\alpha}=\operatorname*{arg\,min}_{\nu\in\arg\max V_{0}}D_{KL}(\nu\|\phi)$
+which, ex-post, is both unsurprising (entropic regularization naturally leads to selecting for the maximum entropy choice) and a reasonable choice.
+Therefore, what we propose is not to find $\nu_{\alpha}^{*}$ for a sequence of $\alpha$’s, which would be cumbersome, but to simply select the minimum divergence marginal.
+The trick is to write $V_{\alpha}$ in terms of the dual.
+Instead of optimizing the dual over $a,b$, we plug the Schrödinger equation into the dual and optimize only over $b$.
+Moreover, since we get one normalization, we restrict $\int b\>d\mu=0$.
+Further, note that if $a$ satisfies the Schrödinger equation, then
+$\int e^{u-a-b}\>d\phi^{\alpha}d\nu^{1-\alpha}\>d\mu-1=0$
+so from the dual, we get
+$V_{\alpha}(\nu)$ $=\inf_{b\ s.t.\int b\ d\mu=0}\int\log\left(\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\right)+\log\left(\int e^{u-b}d\mu\right)\>d\nu$
+$=-\alpha D_{KL}(\nu\|\phi)+\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)d\nu$
+The thing to note here is that $b_{\nu}$, which minimizes $\int\log\left(\int e^{u-b}d\mu\right)\>d\nu$ subject to $\int b\>d\mu=0$, does not depend on $\alpha$.
+So, $\alpha\mapsto V_{\alpha}(\nu)$ is simply an affine function with slope $D_{KL}(\nu\|\phi)$ and $y$-intercept $\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)\ d\nu$.
+Then, define
+$\alpha\mapsto\sup_{\nu}V_{\alpha}(\nu)$
+to be the upper envelope.
+The pointwise supremum of a family of affine functions is convex (and trivially, continuous).
+As $\alpha\to 0$, $D_{KL}(\nu^{*}_{\alpha}\|\phi)$ increases, and therefore, if $\lim_{\alpha\to 0}\nu^{*}_{\alpha}$ exists and is in $\arg\max V_{0}$, it must be the element in the argmax which minimizes $D_{KL}(\nu\|\phi)$.
+In general, for all $\alpha>0$, $D_{KL}(\nu_{\alpha}^{*}\|\phi)\geq D_{KL}(\nu_{0}\|\phi)$ for all $\nu_{0}\in\arg\max V_{0}$.
+Lemma 18.
+$\arg\max V_{\alpha}$ is convex for $\alpha\in[0,1)$
 Proof.
-By assumption, $u$ and $\log\frac{d\nu}{d\phi}$ are totally bounded. This then follows from [6, Theorem 6.15]. ∎
+Two observations: $x\mapsto x^{1-\alpha}$ and $x\mapsto x^{\alpha}$ are concave, and $x\mapsto 1/x$ is convex, for $x\in(0,1)$. This means that for $\nu,\gamma\in\Delta X$,
+$\int\frac{e^{u(\xi,\theta)}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)$
+$\leq\int\frac{e^{u(\xi,\theta)}}{\beta\int e^{u(x,\theta)}\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\ d\phi(x)+(1-\beta)\int e^{u(x,\theta)}\left(\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)$
+$\leq\beta\int\frac{e^{u(\xi,\theta)}}{\int e^{u(x,\theta)}\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)+(1-\beta)\int\frac{e^{u(\xi,\theta)}}{\int e^{u(x,\theta)}\left(\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\>d\mu(\theta)$
+$\leq\beta\left(\frac{d\nu}{d\phi}\right)^{\alpha}+(1-\beta)\left(\frac{d\gamma}{d\phi}\right)^{\alpha}$
+$\leq\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{\alpha}$
+Thus,
+$\int\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\ d\mu(\theta)-1\leq 0$
+However,
+$\int\underbrace{\left[\int\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\ d\mu(\theta)-1\right]}_{\leq 0}\ d(\beta\nu+(1-\beta)\gamma)(\xi)=0$
+which means
+$\int\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}\ d\mu(\theta)=1$
+almost surely. Therefore,
+$P(\cdot|\theta):=\int_{\cdot}\frac{e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{-\alpha}}{\int e^{u(x,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(x)}d(\beta\nu+(1-\beta)\gamma)(\xi)$
+is a valid kernel, insofar as $P_{\xi}=\beta\nu+(1-\beta)\gamma$. Now, suppose $\nu,\gamma\in\arg\max V_{0}$. We get
+$V(\beta\nu+(1-\beta)\gamma)$ $\geq U(P)$
+$=\int u(\xi,\theta)-\alpha\log\left(\frac{d(\beta\nu+(1-\beta)\gamma)(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi|\theta)}{d(\beta\nu+(1-\beta)\gamma)(\xi)}\right)\>dP$
+$=\int\log\left(\int e^{u(\xi,\theta)}\left(\beta\frac{d\nu}{d\phi}+(1-\beta)\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(\xi)\right)\>d\mu(\theta)$
+$\geq\int\beta\log\left(\int e^{u(\xi,\theta)}\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\ d\phi(\xi)\right)+(1-\beta)\log\left(\int e^{u(\xi,\theta)}\left(\frac{d\gamma}{d\phi}\right)^{1-\alpha}\ d\phi(\xi)\right)\>d\mu(\theta)$
+$=\beta V(\nu)+(1-\beta)V(\gamma)$
+One final point to make is that the maximum entropy selector tends to select for non-sparse consideration sets – that is, if there are two consideration sets, their union is also a consideration set, and the maximum entropy selector will prefer to select for the union.
+Depending on the context, either maximum entropy or consideration set sparsity can be reasonable descriptions of weak human preferences, but they are inherently contradictory, so if part of the impetus for using a rational inattention model is to induce sparse consideration sets, then using a maximum entropy selector would be counterproductive in spirit.
+And it is no more inherent to use maximum entropy than it is to use minimum cardinality to select among possible outcomes.
+Example. Consider a state-action model with a state space
+$\Theta=\{1,2,3,4\}$ $\mu(\theta)=1/4$
+and an action space
+$A=\bigg\{\{1,2\},\{1,3\},\{1,4\},\{2,3\},\{2,4\},\{3,4\}\bigg\}$
+and a utility
+$u(a,\theta)=\textbf{1}\{\theta\in a\}$
+That is, the DM chooses a pair of states, and gets utility $1$ iff the drawn state corresponds to one in the chosen pair.
+It is easy to see that any pair of actions whose disjoint union is $\Theta$ is a valid consideration set.
+One need only check the necessary and sufficient condition provided by [caplin2019rational].
+Any convex combination of optimal marginals is also an optimal marginal given that $V$ is concave, meaning a union of two optimal consideration sets is an optimal consideration set.
+So, $A$ itself is a valid consideration set.
+What is not a valid (optimal) consideration set, however, is something like $\bigg\{\{1,2\},\{2,3\},\{3,4\}\bigg\}$.
+[^7] It is easy to show that if $\nu(2,3)>0$, then $\sum_{\theta}\frac{e^{u((2,3),\theta)}}{\sum_{a}\nu(a)e^{u(a,\theta)}}\mu(\theta)>1$ The idea is that if the consideration set was $\bigg\{\{1,2\},\{2,3\},\{3,4\}\bigg\}$ then receiving the action recommendation signal $\{1,2\}$ provides additional but not-utility-relevant information about whether the true state is $1$.
+In an optimal information policy, the action recommendation $(i,j)\in A$ should not differentiate (reveal unnecessary information) about whether the true state is $i$ or $j$, because once the DM chooses $(i,j)$, such information would be irrelevant.
+To see this in an extreme case, suppose $\nu(2,3)=0.99$.
+Then, receiving the action recommendation $(1,2)$ would strongly indicate that the true state is $1$.
+The point is that there are sparse optimal consideration sets like $\bigg\{\{1,2\},\{3,4\}\bigg\}$ and non-sparse ones, like $A$.
+The maximum entropy selector, however, will choose the non-sparse one, assuming that $\phi(a)=1/6$.
+And indeed, it will be the case $\nu^{*}_{0}(a)=1/6$.
+Fixed-Point Iteration
+Let $X$ be discrete.
+Let $\delta_{x}$ denote the Dirac measure at $x\in X$ and let $\partial_{x}f(\nu)$ be the Gateaux derivative in the direction of $\delta_{x}$.
+Again, let $g=d\nu/d\phi$, generically.
+We abuse notation by using $g$ and $\nu$ interchangeably: we use $f(g)$ to mean $f(\nu)$, $\partial_{x}f(g)$ to mean $\partial_{x}f(\nu)$, $T\nu$ to mean $Tg$, etc.
+It should be clear in each instance whether we are referring to the density or the measure.
+Lemma 19.
+Let $k\geq 1-\alpha$ and $n\geq 0$. Then,
+$\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{k}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{n}\>d\phi(\xi)\right)\>d\mu(\theta)$
+$\leq(1-\alpha)f(Tg)+\alpha\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\>d\phi(\xi)\right)\>d\mu(\theta)$
+with equality only if for $\mu$-a.e. $\theta$,
+$e^{u(\xi,\theta)}[Tg](\xi)^{1-\alpha}\text{ is linearly dependent on }e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\text{ in }L^{1}(\phi)$
+Proof.
+We exploit the fact that
+$\partial_{\xi}f(g)=(1-\alpha)\left(\frac{Tg(x)}{g(x)}-1\right)$
+Re-arrange to get
+$g(x)=\frac{Tg(x)}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}$
+We can re-write $g(\xi)$ as
+$g(\xi)^{k}=g(\xi)^{k-(1-\alpha)^{2}}\left(\frac{Tg(\xi)}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{(1-\alpha)^{2}}$
+Plugging this in, we get
+$\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{k}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{n}\>d\phi(\xi)\right)\>d\mu(\theta)$
+$=\int\log\left(\int e^{\alpha u(\xi,\theta)}e^{(1-\alpha)u(\xi,\theta)}[Tg](\xi)^{(1-\alpha)^{2}}g(\xi)^{k-(1-\alpha)^{2}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{n+(1-\alpha)^{2}}\>d\phi(\xi)\right)\>d\mu(\theta)$
+$\leq\int\log\left[\left(\int e^{u(\xi,\theta)}[Tg](\xi)^{1-\alpha}d\phi(\xi)\right)^{1-\alpha}\left(\int e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\>d\phi(\xi)\right)^{\alpha}\right]\>d\mu(\theta)$
+$=(1-\alpha)f(Tg)+\alpha\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{\frac{k-(1-\alpha)^{2}}{\alpha}}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{\frac{n+(1-\alpha)^{2}}{\alpha}}\>d\phi(\xi)\right)\>d\mu(\theta)$
+via Holder’s inequality. ∎
+Starting with $k=1-\alpha$, we see that $\frac{1-\alpha-(1-\alpha)^{2}}{\alpha}=(1-\alpha)\left(\frac{1-(1-\alpha)}{\alpha}\right)=1-\alpha=k$.
+Starting from $n=0$, by iteratively adding $(1-\alpha)^{2}$ and dividing by $\alpha$, we get convergence to
+$(1-\alpha)^{2}\left(\frac{1}{\alpha}+\frac{1}{\alpha^{2}}+\frac{1}{\alpha^{3}}+\cdots\right)=(1-\alpha)^{2}\left(\frac{1}{1-\alpha}-1\right)$
+Thus,
+$f(g)$ $=\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{1-\alpha}\>d\phi(\xi)\right)\ d\mu(\theta)$
+$\leq(1-\alpha)f(Tg)+\alpha\int\log\left(\int e^{u(\xi,\theta)}g(\xi)^{1-\alpha}\left(\frac{1}{1+\frac{\partial_{\xi}f(g)}{1-\alpha}}\right)^{(1-\alpha^{2})\left(\frac{1}{\alpha}\right)}\ d\phi(\xi)\right)\ d\mu(\theta)$
+$\ \vdots$
+$\leq(1-\alpha)\left(1+\alpha+\cdots+\alpha^{k}\right)f(Tg)+\alpha^{k+1}\int\log\left(\int\frac{e^{u(\xi,\theta)}g(\xi)^{1-\alpha}}{\left(1+\frac{\partial_{\xi}f(g)}{1-\alpha}\right)^{(1-\alpha)^{2}\left(\frac{1}{\alpha}+\cdots+\frac{1}{\alpha^{k+1}}\right)}}\ d\phi(\xi)\right)\>d\mu(\theta)$
+$\ \vdots$
+$\leq f(Tg)$
+Note that the inequalities, which come from Holder, are usually strict.
+In order for the inequality to not be strict, $e^{u(\xi,\theta)}[Tg](\xi)^{1-\alpha}$ must be $\mu$-almost surely linearly dependent on $\frac{e^{u(\xi,\theta)}g(\xi)^{1-\alpha}}{\left(1+\frac{\partial_{\xi}f(g)}{1-\alpha}\right)^{(1-\alpha)^{2}\left(\frac{1}{\alpha}+\cdots+\frac{1}{\alpha^{k+1}}\right)}}$.
+Even if this holds for some $k$, it won’t hold for $k+1$ unless $\partial_{\xi}f(g)=0$, either of which suffice for optimality.